### v1

# Directional Convergence, Benign Overfitting of Gradient Descent in leaky ReLU two-layer Neural Networks

Abstract
In this paper, we prove directional convergence of network parameters of fixed width leaky ReLU two-layer neural networks optimized by gradient descent with exponential loss, which was previously only known for gradient flow. By a careful analysis of the convergent direction, we establish sufficient conditions of benign overfitting and discover a new phase transition in the test error bound. All of these results hold beyond the nearly orthogonal data setting which was studied in prior works. As an application, we demonstrate that benign overfitting occurs with high probability in sub-Gaussian mixture models.
1 Introduction
The practical success of deep learning has revealed surprising theoretical phenomena. One such phenomenon, which has inspired intense theoretical research, is benign overfitting, where over-parametrized neural network models can achieve arbitrarily small test errors while perfectly interpolating training data—even in the presence of label noise (Zhang et al. (2017); Belkin et al. (2019)). Benign overfitting has attracted broad attention since it seems to conflict with the classical statistical understanding that there should be a trade-off between fitting the training data and generalization on the test data. This led to intense theoretical research which has revealed that benign overfitting also occurs in several classical statistical models, such as linear regression (Bartlett et al. (2020), Muthukumar et al. (2020), Mei and Montanari (2022), Hastie et al. (2022)), ridge regression (Tsigler and Bartlett (2024)), binary linear classification (Chatterji and Long (2021); Wang and Thrampoulidis (2022); Cao et al. (2021); Hashimoto et al. (2025)), and kernel-based estimators (Liang and Rakhlin (2020), Liang et al. (2020)), to name a few. However, our understanding of the phenomenon for neural networks remains limited.
It is widely believed that the implicit bias, which results from implicit regularization of gradient based optimization, is the key to understand generalization performance in the over-parametrized regime. In fact, for binary linear classification, it is known that linear classifiers trained under gradient descent with a loss function with a tight exponential tail, e.g. the exponential loss and logistic loss, converge in direction to the maximum margin classifier (Soudry et al. (2018)). Benign overfitting in binary linear classification has been studied thoroughly by analyzing the convergent direction (Wang and Thrampoulidis (2022); Cao et al. (2021); Hashimoto et al. (2025)).
In this paper, we investigate benign overfitting in two-layer leaky ReLU neural networks with fixed width for binary classification whose network parameters are optimized by gradient descent with exponential loss.
Our main results are summarized as follows:
  1. 1.
    Directional Convergence: we show that the network parameter converges in direction if the parameter is optimized by gradient descent with exponential loss. In the existing literature, the directional convergence was proven only for gradient flow by Ji and Telgarsky (2020). For gradient descent, only weaker forms of implicit bias were shown (Lyu and Li (2020); Frei et al. (2023b); Kou et al. (2023b)). Moreover, the analysis in Frei et al. (2023b); Kou et al. (2023b) is limited to the rather specialized setting of nearly orthogonal data, i.e. the predictors $\{\boldsymbol{x}_{i}\}_{i=1,\ldots,n}\subset\mathbb{R}^{p}$ satisfy $\|\boldsymbol{x}_{i}\|^{2}\gg\max_{i\neq k}|\langle\boldsymbol{x}_{i},\boldsymbol{x}_{k}\rangle|$.
  2. 2.
    Benign Overfitting: as an application of the directional convergence, we establish sufficient conditions for benign overfitting by analyzing the property of convergent direction. Due to the difficulty in establishing directional convergence, prior works for leaky ReLU networks were limited to either gradient flow (Frei et al. (2023a)), smoothed approximation (Frei et al. (2022)), or nearly orthogonal data regime (Xu and Gu (2023)). Taking advantage of a careful analysis of the convergent direction, our work establishes sufficient conditions for benign overfitting beyond previously analyzed regimes, which reveals a new phase transition in the test error bound.
  3. 3.
    Beyond NTK: both the directional convergence and benign overfitting are shown for fixed width two-layer neural networks, thus our result is beyond the so called neural tangent kernel (NTK) regime or the lazy training regime (Jacot et al. (2018); Arora et al. (2019)).
2 Related Work
Implicit bias in neural networks for binary classification.
Here, we focus on works on leaky ReLU homogeneous neural networks which are closely related to ours. We refer readers to Vardi (2022) for a comprehensive survey on implicit bias in neural networks.
Most prior works on the implicit bias of (leaky) ReLU neural networks for binary classification have considered gradient flow. For gradient flow, Lyu and Li (2020) showed that any limit point of the network parameter of homogeneous neural networks trained with exponential type loss is a KKT point of a constrained optimization problem and that margin maximization occurs. Ji and Telgarsky (2020) further established directional convergence of the parameter of homogeneous neural networks trained with the exponential type loss. Taking advantage of the directional convergence of gradient flow, Lyu et al. (2021); Bui Thi Mai and Lampert (2021); Sarussi et al. (2021); Frei et al. (2023b); Min et al. (2024) obtained detailed characterizations of the convergent direction of the parameters of (leaky) ReLU two-layer neural networks under further assumptions on on the training data.
In contrast to gradient flow, directional convergence of network parameters was not proven for gradient descent with exponential type loss, even under strong assumptions. As a result, only weaker forms of implicit bias have been obtained for gradient descent. Lyu and Li (2020) showed that margin maximization also occurs for the network parameters of a homogeneous network trained by gradient descent, but their assumptions include smoothness of the network, and hence rule out (leaky) ReLU networks. Frei et al. (2023b) showed that the stable rank of network parameters stays bounded by a constant for two-layer neural networks, but their result is limited to the smoothed approximation of leaky ReLU. Kou et al. (2023b) proved similar results for (leaky) ReLU two-layer neural networks assuming that the number of positive and negative neurons are equal. Moreover, Frei et al. (2023b); Kou et al. (2023b) are limited to the rather specialized scenario of nearly orthogonal data.
Benign overfitting in neural networks for binary classification.
As noted earlier, benign overfitting in neural networks remains poorly understood. In fact, very few theoretical results are established beyond the NTK regime. Cao et al. (2022) studied benign overfitting for two-layer convolutional neural networks with polynomial ReLU activation ($\mathrm{ReLU}^{q},q>2$). Their results were strengthened by Kou et al. (2023a) by relaxing the polynomial ReLU assumption. Both Cao et al. (2022) and Kou et al. (2023a) are not applicable to our setting since our work considers fully connected neural networks.
More closely related work is Frei et al. (2022), which studied benign overfitting in smoothed leaky ReLU two-layer neural networks on mixture data. Their work was extended by Xu and Gu (2023) to non-smooth leaky ReLU like activation. While Xu and Gu (2023)’s work is not limited to (leaky) ReLU, their work is limited to the nearly orthogonal data regime and only holds for isotropic mixture with bounded log-Sobolev constant, which is a stronger assumption than sub-Gaussian mixture. Our work establishes benign overfitting beyond the nearly orthogonal data regime and can be applied to a wider class of mixtures than prior works.
A concise summary of related prior works is given in Table 1.
Table 1: Summary of Existing Studies of leaky ReLU two-layer neural networks
Directional Convergence Benign Overfitting
Nearly Orthogonal Data Other regime Nearly Orthogonal Data Other regime
Gradient Flow Ji and Telgarsky (2020) Frei et al. (2023a) NA
Gradient Descent NA* NA Frei et al. (2022)**, NA
Xu and Gu (2023)
* Some weaker forms of implicit bias were studied by Frei et al. (2023b); Kou et al. (2023b).
** Frei et al. (2022) studied smoothed approximation of leaky ReLU instead.
3 Preliminaries
We consider two-layer neural networks
$f(\boldsymbol{x};W)=\sum_{j=1}^{m}a_{j}\phi(\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle),$ (1)
where $\boldsymbol{x}\in\mathbb{R}^{p}$, $\boldsymbol{w}_{j}\in\mathbb{R}^{p}$, $W=(\boldsymbol{w}_{1},\boldsymbol{w}_{2},\ldots,\boldsymbol{w}_{m})\in\mathbb{R}^{p\times m}$, $|a_{j}|=\pm\tfrac{1}{\sqrt{m}}$ and $\phi$ is $\gamma-$leaky ReLU activation, i.e. $\phi(x)=\max\{x,\gamma x\}$ for $\gamma\in(0,1)$. Letting $\zeta(x)=1$ if $x\geq 0$ and $\zeta(x)=\gamma$ if $x<0$, we may conveniently write $\phi(x)=\zeta(x)x$.
Letting $J=\{1,2,\dots,m\}$, $J_{+}=\{j\in J:a_{j}=\tfrac{1}{\sqrt{m}}\}$, and $J_{-}=J\setminus J_{+}$, we can rewrite the network (1) as
$f(\boldsymbol{x};W)=\sum_{j\in J_{+}}\frac{1}{\sqrt{m}}\phi(\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle)-\sum_{j\in J_{-}}\frac{1}{\sqrt{m}}\phi(\langle\boldsymbol{x},\boldsymbol{w}_{j}\rangle).$
We consider the exponential loss $\ell(u)=\exp(-u)$ and define the empirical loss of the dataset $\{(\boldsymbol{x}_{i},y_{i})\}_{i\in I}$, where $I:=\{1,2,\ldots,n\}$, by
$\mathcal{L}(W)=\sum_{i=1}^{n}\ell(y_{i}f(\boldsymbol{x}_{i};W))=\sum_{i=1}^{n}\exp(-y_{i}f(\boldsymbol{x}_{i};W)).$
We train the first layer weights $\boldsymbol{w}_{j}$ by gradient descent initialized at $W^{(0)}$, i.e.
$W^{(t+1)}=W^{(t)}-\alpha\nabla_{W}\mathcal{L}(W^{(t)}),\quad t=0,1,2,\ldots,$ (2)
where $\alpha>0$ is a fixed constant and we assume $\phi^{\prime}(0)$ is any number in $[\gamma,1]$. The second layer weights $a_{j}$ are fixed throughout the training. Fixing the second layer weights $a_{j}\in\{\pm 1/\sqrt{m}\}$ and only optimizing the first layer weights $\boldsymbol{w}_{j}$ is standard in the existing literature, see e.g. Arora et al. (2019); Cao et al. (2022); Frei et al. (2022); Xu and Gu (2023); Kou et al. (2023b). We do not consider random initialization of $W^{(0)}$ since only the magnitude of $W^{(0)}$ plays a role in our analysis.
Assume that we observe i.i.d. copies $\{(\boldsymbol{x}_{i},y_{i})\}_{i=1,\ldots,n}$ from a mixture distribution on $\mathbb{R}^{p}\times\{\pm 1\}$ which is defined as follows
  1. (M)
    The observations consist of $n$ i.i.d copies $(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n$ of the pair $(\boldsymbol{x},y)$. Here for a random variable $y\in\{-1,1\}$ satisfying $P(y=1)=P(y=-1)=1/2$, a random vector $\boldsymbol{z}\in\mathbb{R}^{p}$ independent of $y$, and a deterministic $\boldsymbol{\mu}\in\mathbb{R}^{p}$, we have
    $\boldsymbol{x}\;=\;y\boldsymbol{\mu}+\boldsymbol{z}.$ (3)
The sub-Gaussian norm of a random variable $X$, say $\|X\|_{\psi_{2}}$, is defined as
$\|X\|_{\psi_{2}}:=\inf\{t>0:\,\mathbb{E}\exp(X^{2}/t^{2})\leq 2\}.$
The sub-Gaussian norm of a random vector $\boldsymbol{X}\in\mathbb{R}^{p}$ is denoted by $\|\boldsymbol{X}\|_{\psi_{2}}$ and defined as
$\|\boldsymbol{X}\|_{\psi_{2}}:=\sup_{\boldsymbol{v}\in S^{p-1}}\|\langle\boldsymbol{v},\boldsymbol{X}\rangle\|_{\psi_{2}}.$
As a special case of the model (M), we consider sub-Gaussian mixture models:
  1. (sG)
    Suppose in model (M) $\boldsymbol{z}=\Sigma^{\frac{1}{2}}\boldsymbol{\xi}$, where $\boldsymbol{\xi}\in\mathbb{R}^{p}$ has independent entries $\xi_{j}$ that have mean zero, unit variance, and satisfy $\|\xi_{j}\|_{\psi_{2}}\leq L$ for all $j=1,\ldots,p$.
The mixture models (M) and (sG) have been studied for linear classifiers by Chatterji and Long (2021); Wang and Thrampoulidis (2022); Hashimoto et al. (2025) and for two-layer neural networks with $\Sigma=I_{p}$ by Frei et al. (2022); Xu and Gu (2023).
We denote by $\|A\|$ and $\|A\|_{F}$ the spectral norm and the Frobenius norm of a matrix $A$, respectively.
4 Directional Convergence
We start from our main results on directional convergence of leaky ReLU two-layer neural networks.
The key to show the directional convergence of the gradient descent iterates (2) is activation of neurons. We say that $j$-th neuron $\phi(\langle\cdot,\boldsymbol{w}_{j}\rangle)$ is activated if $a_{j}y_{i}\phi(\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle)>0$ holds for any $i\in I$, which can be equivalently written as $a_{j}y_{i}\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle>0$ for any $i\in I$. We will later establish conditions under which $a_{j}y_{i}\langle\boldsymbol{x}_{i},w_{j}^{(t)}\rangle$ monotonically increases as the gradient descent progresses.
Our first result shows that if all the neurons stay activated after some step of gradient descent (2), directional convergence occurs towards the maximum margin vector $\hat{W}$ in a transformed sample space, i.e. $\hat{W}$ is the solution to the following optimization problem:
$\text{Minimize}\quad\|\boldsymbol{w}\|^{2}\quad\text{s.t.}\quad\langle y_{i}\boldsymbol{\tilde{x}}_{i},\boldsymbol{w}\rangle\geq 1\quad\text{for all $i\in I$, where $\boldsymbol{w}=\mathrm{vec}(W)$},$ (4)
where $\boldsymbol{\tilde{x}}_{i}^{\top}=(\boldsymbol{\tilde{x}}_{i,1}^{\top},\boldsymbol{\tilde{x}}_{i,2}^{\top},\ldots,\boldsymbol{\tilde{x}}_{i,m}^{\top})\in\mathbb{R}^{1\times mp}$ with $\boldsymbol{\tilde{x}}_{i,j}=a_{j}\zeta(a_{j}y_{i})\boldsymbol{x}_{i}$. Letting $\sigma_{max}(\tilde{X})$ be the maximum singular value of $\tilde{X}\in\mathbb{R}^{n\times mp}$ whose rows consist of $\boldsymbol{\tilde{x}}_{i}^{\top}$’s, we have
Theorem 4.1.
Suppose all the neurons are activated for $t\geq T$ and $\alpha\sigma_{max}^{2}(\tilde{X})$ <2. Then, the gradient descent iterate (2) satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$ and $W^{(t)}$ converges in direction to $\hat{W}$.
Furthermore, the convergent direction $\{\boldsymbol{\hat{w}}_{j}\}_{j\in J}$ can also be given by $\boldsymbol{\hat{w}}_{j}=\boldsymbol{w}_{+}$ for $j\in J_{+}$ and $\boldsymbol{\hat{w}}_{j}=\boldsymbol{w}_{-}$ for $j\in J_{-}$, which are the solution to the following optimization problem:
$\text{Minimize}\quad|J_{+}|\|\boldsymbol{w}_{+}\|^{2}+|J_{-}|\|\boldsymbol{w}_{-}\|^{2}$ (5)
s.t. $\quad\frac{|J_{+}|}{\sqrt{m}}\langle\boldsymbol{x}_{i},\boldsymbol{w}_{+}\rangle-\frac{\gamma|J_{-}|}{\sqrt{m}}\langle\boldsymbol{x}_{i},\boldsymbol{w}_{-}\rangle\geq 1,\forall i\in I_{+},$
$\quad\frac{|J_{-}|}{\sqrt{m}}\langle\boldsymbol{x}_{i},\boldsymbol{w}_{-}\rangle-\frac{\gamma|J_{+}|}{\sqrt{m}}\langle\boldsymbol{x}_{i},\boldsymbol{w}_{+}\rangle\geq 1,\forall i\in I_{-}.$
Lastly, the network $f(\cdot;\hat{W})$ has the linear decision boundary defined by $\boldsymbol{\bar{w}}:=\frac{|J_{+}|}{\sqrt{m}}\boldsymbol{w}_{+}-\frac{|J_{-}|}{\sqrt{m}}\boldsymbol{w}_{-}$, i.e. $\mathrm{sign}(f(\boldsymbol{x}:\hat{W}))=\mathrm{sign}(\langle\boldsymbol{x},\boldsymbol{\bar{w}}\rangle)$ for any $\boldsymbol{x}\in\mathbb{R}^{p}$.
Previously, the characterizations of the convergent direction in Theorem 4.1 were only obtained for gradient flow trained on nearly orthogonal data by Frei et al. (2023b). The proof of Theorem 4.1 is given in Appendix A.1.
To demonstrate that Theorem 4.1 can guarantee directional convergence of the network parameter $W^{(t)}$ beyond the nearly orthogonal regime, we verify the conditions of the theorem in model (M). Specifically, we find conditions which ensure that the neurons stay activated and $\alpha\sigma_{max}^{2}(\tilde{X})<2$.
In the following, we let $R_{max}(\boldsymbol{z}):=\max_{i}\|\boldsymbol{z}_{i}\|$ and $R_{min}(\boldsymbol{z}):=\min_{i}\|\boldsymbol{z}_{i}\|$, and define events that capture deterministic conditions the training data need to satisfy for the directional convergence:
$E(\theta_{1},\theta_{2})=\Big{\{}\max_{i\neq k}|\langle\tfrac{\boldsymbol{z}_{i}}{\|\boldsymbol{z}_{i}\|},\tfrac{\boldsymbol{z}_{k}}{\|\boldsymbol{z}_{k}\|}\rangle|\leq\theta_{1}\quad\text{and}\quad\max_{i}|\langle\tfrac{\boldsymbol{z}_{i}}{\|\boldsymbol{z}_{i}\|},\tfrac{\boldsymbol{\mu}}{\|\boldsymbol{\mu}\|}\rangle|\leq\theta_{2}\Big{\}},$ (6)
for some $\theta_{1},\theta_{2}\in[0,1]$ (take $\theta_{1}=0$ if $\boldsymbol{\mu}=0$ or $R_{max}(\boldsymbol{z})=0$ and take $\theta_{2}=0$ if $\boldsymbol{\mu}=0$ or $R_{max}(\boldsymbol{z})=0$). Event $E$ allows us to control $\|\boldsymbol{x}_{i}\|$ and $\langle\boldsymbol{x}_{i},\boldsymbol{x}_{k}\rangle,i\neq k$ in terms of $\|\boldsymbol{\mu}\|,R_{min}(\boldsymbol{z}),R_{max}(\boldsymbol{z}),\theta_{1}$, and $\theta_{2}$.
In addition, we denote by $\sigma:=\max_{j}\|\boldsymbol{w}_{j}\|$ the size of initialization. It is also convenient to let $\rho:=\sigma\sqrt{m(1+\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}}$ to make the assumptions of our main results concise.
We present the sufficient conditions of directional convergence for the following two cases:
Case 1: $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$.
In this case, we make the following assumptions:
Assumption 4.2 (Small Initialization).
$\alpha>\frac{\rho\exp(\rho)}{\gamma(1-\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{min}^{2}(\boldsymbol{z})\}}$
Assumption 4.3 (Small Step Size).
$\alpha(n\|\boldsymbol{\mu}\|^{2}+R_{max}(\boldsymbol{z})^{2})<1.$
The intuition behind these assumption is as follows: It can be shown that $\langle y_{i}\boldsymbol{x}_{i},y_{k}\boldsymbol{x}_{k}\rangle\geq 0$ for any $i\neq k$ under $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$. Assumption 4.2 ensures the step size sufficiently larger than initialization so that neuron activation occurs immediately after one step of (2). Then, Assumption 4.3 ensures $\alpha\tilde{\sigma}_{max}^{2}(\tilde{X})<2$ so that the directional convergence holds.
Theorem 4.4 (Directional Convergence, Case 1).
Suppose event $E$ holds with Assumptions 4.2 and 4.3 with $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ and $\theta_{2}<1$. Then, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, and converges in the direction given by Theorem 4.1.
We note that there is no upper bound to the signal $\|\boldsymbol{\mu}\|$. Therefore, Theorem 4.4 allows non-nearly-orthogonal data. The proof of Theorem 4.4 is given in Appendix A.2.
Case 2: Weak Signal Regime.
To deal with the case when the signal is not strong enough so that $\langle y_{k}\boldsymbol{x}_{k},y_{i}\boldsymbol{x}_{i}\rangle,k\neq i$ could be negative, we consider the following assumptions:
Assumption 4.5 (Near Orthogonality).
$2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})\leq\frac{\varepsilon_{1}\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})}{n\exp(2\rho)}\quad\text{for some $\varepsilon_{1}\in[0,1]$}.$
Assumption 4.6 (Small Initialization).
$\alpha\geq\frac{\rho\exp(\rho)}{\varepsilon_{2}\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})}\quad\text{for some $\varepsilon_{2}\in(0,1]$}.$
Assumption 4.7 (Weak Signal).
$n\|\boldsymbol{\mu}\|^{2}\leq\varepsilon_{3}R_{min}^{2}(\boldsymbol{z})\quad\text{for some $\varepsilon_{3}\in[0,1)$}.$
Assumption 4.8 (Small Step Size).
$16\alpha R_{max}^{2}(\boldsymbol{z})\exp(\rho)<$ 1 .
Assumption 4.5 and 4.7 allow us to carefully control $\langle y_{k}\boldsymbol{x}_{k},y_{i}\boldsymbol{x}_{i}\rangle$. Small $\varepsilon_{1},\varepsilon_{3}$ ensure that the training data are nearly orthogonal, and small $\varepsilon_{2}$ requires sufficiently small initialization relatively to step size. By taking $\boldsymbol{\mu}=0$, this case includes the nearly orthogonal data regime studied in prior works (Frei et al. (2023b); Kou et al. (2023b)). The other assumptions are analogous to Case 1.
With these assumptions, we have
Theorem 4.9 (Directional Convergence for Weak Signal Regime, Case 2).
Suppose event $E$ holds under Assumption 4.5, 4.6, 4.7, and 4.8 with $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$, $C_{w}\varepsilon_{i}\leq\tfrac{1}{2},i=1,3$, and $\theta_{2}\leq\tfrac{1}{2}$, where $C_{w}$ is a constant defined by
$C_{w}:=\frac{24R_{max}^{2}(\boldsymbol{z})}{\gamma^{2}R_{min}^{2}(\boldsymbol{z})}\exp(1).$
Then, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, and converges in the direction given by Theorem 4.1.
The proof of Theorem 4.9 is given in Appendix A.3.
We note that Case 1 was not studied in prior works while Case 2 essentially corresponds to the near orthogonal regime studied in Frei et al. (2023b); Kou et al. (2023b). Therefore, theorems in this section not only strengthen existing results but also establish directional convergence in a regime not studied in the existing literature.
5 Classification Error of the Convergent Direction
In this section, we discuss the classification performance of the convergent direction $\hat{W}$ of (2) whose characterization given by Theorem 4.1. We establish a classification error bound by obtaining a closed form expression of $\boldsymbol{\hat{w}}=\mathrm{vec}(\hat{W})$, which is $\boldsymbol{\hat{w}}=\tilde{X}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$, where $\boldsymbol{y}=(y_{1},\ldots,y_{n})^{\top}\in\mathbb{R}^{n}$.
Near orthogonality of $\boldsymbol{z}_{i}$’s to each other and to $\boldsymbol{\mu}$ is again the key to obtain the closed form expression and to establish an error bound. Letting $Z=(\boldsymbol{z}_{1},\ldots,\boldsymbol{z}_{n})^{\top}\in\mathbb{R}^{n\times p}$, $n_{+}=|\{i\in I:y_{i}=1\}|$, and $n_{-}=n-n_{+}$, the near orthogonality condition is characterized by the parameters from the following events which are defined for $R\geq 0,\tilde{\varepsilon}_{1}\geq 0,\tilde{\varepsilon}_{2}\geq 0$, and $\tilde{\varepsilon}_{3}\in[0,1]$:
$\tilde{E}_{1}(\tilde{\varepsilon}_{1},R)$ :=$:=\left\{\|ZZ^{\top}-RI_{n}\|\leq\tfrac{\tilde{\varepsilon}_{1}}{3}R\right\},$ (7)
$\tilde{E}_{2}(\tilde{\varepsilon}_{2},R)$ :=$:=\left\{\|Z\boldsymbol{\mu}\|\leq\tfrac{\tilde{\varepsilon}_{2}}{3}R\right\},$ (8)
$\tilde{E}_{3}(\tilde{\varepsilon}_{3})$ :=$:=\left\{|n_{+}-\tfrac{n}{2}|\leq\tilde{\varepsilon}_{3}\tfrac{n}{2}\right\}.$ (9)
Event $\tilde{E}_{1}$ captures near orthogonality of $\boldsymbol{z}_{i}$’s to each other and uniformity of their norms, and event $\tilde{E}_{2}$ captures near orthogonality of $\boldsymbol{z}_{i}$’s to $\boldsymbol{\mu}$. Although we could use the same parameters as in the analysis of directional convergence, these events $\tilde{E}_{1},\tilde{E}_{2}$ makes the analysis of $\hat{W}$ simpler.
With $\tilde{\varepsilon}:=\max\{\tilde{\varepsilon}_{1},\sqrt{n}\tilde{\varepsilon}_{2}\}$, $q_{+}=|J_{+}|/m$, $q_{-}=|J_{-}|/m$ , and $q_{\gamma}=\min\{q_{+}+\gamma^{2}q_{-},q_{-}+\gamma^{2}q_{+}\}$, we can now present our main result on the classification performance of $\hat{W}$:
Theorem 5.1.
Suppose event $\bigcap_{i=1,2,3}\tilde{E}_{i}$ holds with $\tilde{\varepsilon}\leq\tfrac{q_{\gamma}}{2}$, $\tilde{\varepsilon}_{3}\leq\tfrac{1}{2}$, and further assume one of the following condition holds for a sufficiently large constant $C=C(\gamma,q_{+})$:
  1. (i)
    $Cn\|\boldsymbol{\mu}\|^{2}\leq R$, $C\tilde{\varepsilon}\leq n^{-\frac{1}{2}}$, and $C\tilde{\varepsilon}_{2}\leq\frac{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}{R}$,
  2. (ii)
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq CR$, $C\tilde{\varepsilon}\leq\frac{R}{n^{\frac{3}{2}}\|\boldsymbol{\mu}\|^{2}}$, and $C\tilde{\varepsilon}_{2}\leq n^{-\frac{1}{2}}$.
Then, we have
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(x;\hat{W})<0)\leq\exp\left(-c\frac{n\|\boldsymbol{\mu}\|^{4}}{\|\boldsymbol{z}\|_{\psi_{2}}^{2}\{n\|\boldsymbol{\mu}\|^{2}+R\}}\right),$
where $c$ is a constant which depends on $\gamma$ and $q_{+}$.
The theorem statement here is a concise versions for the purpose of presentation. Detailed version of the theorem with the explicit form of constant $C$ and the proof are given in Theorem B.11, Appendix B.
Theorem 5.1 shows that there is a phase transition between the weak signal regime, i.e. $n\|\boldsymbol{\mu}\|^{2}\lesssim R$ and the strong signal regime, i.e. $n\|\boldsymbol{\mu}\|^{2}\gtrsim R$. The weak signal regime corresponds to the nearly orthogonal data regime studied in prior works under specific distributional settings. Specifically, Frei et al. (2022) assumed $\boldsymbol{z}$ is sub-Gaussian and strongly log-concave and Xu and Gu (2023) assumed $\boldsymbol{z}$ has a bounded log-Sobolev constant, which also is stronger than assuming $\boldsymbol{z}$ is sub-Gaussian. Furthermore, both works only allow for isotropic $\boldsymbol{z}$. Our result not only extends the prior works to the strong signal regime but also establishes sufficient deterministic conditions on the data to guarantee benign overfitting, which allows us to apply the result to a wider distributional setting.
A similar phase transition in the classification error bound was previously established for binary linear classification in Cao et al. (2021); Wang and Thrampoulidis (2022); Hashimoto et al. (2025).
Lastly, we note that the condition $q_{\gamma}>\gamma$ in (iii) is necessary to ensure the equivalence of $\boldsymbol{\hat{w}}$ and the minimum norm least square estimator. In fact, we can prove that the equivalence fails if $q_{\gamma}<\gamma$, which can be viewed as another form of the phase transition in the behavior of the convergent direction. Additional details are provided in Appendix B.
6 Benign Overfitting on sub-Gaussian Mixture
We here present applications of Theorem 4.4, 4.9, and 5.1 to anisotropic sub-Gaussian mixtures (sG). Our main results are given separately for the strong signal and weak signal regimes. For the strong signal regime, we can show
Theorem 6.1 (Strong Signal Regime).
Consider model (sG). Let $\delta\in(0,\tfrac{1}{4})$. Suppose $q_{\gamma}>\gamma$, $n\geq C$, $\alpha\leq\frac{1}{2.4n\|\boldsymbol{\mu}\|^{2}}$, $\sigma\leq 0.2\gamma\alpha\sqrt{\frac{\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)}{m}}$,
${\rm tr}(\Sigma)$ $\geq C\max\Big{\{}n\|\boldsymbol{\mu}\|\|\Sigma\|_{F}^{\frac{1}{2}},n^{5/4}\|\boldsymbol{\mu}\|\|\Sigma\|^{\frac{1}{2}},n^{\frac{5}{4}}\|\boldsymbol{\mu}\|\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|^{\frac{1}{2}},n\frac{\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|^{2}}{\|\boldsymbol{\mu}\|^{2}},\sqrt{n}\|\Sigma\|_{F},n\|\Sigma\|\Big{\}},$
$\|\boldsymbol{\mu}\|^{2}$ $\geq C\max\Big{\{}\frac{{\rm tr}(\Sigma)}{n},\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|,\sqrt{n}\|\Sigma\|_{F},n\|\Sigma\|\Big{\}}$
for a sufficiently large constant $C$ which depends only on $\delta,\gamma,q_{+}$. Then, we have with probably at least $1-3\delta$: the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, converges in direction, and the convergent direction $\hat{W}$ attains
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(x;\hat{W})<0)\leq\exp\left(-\frac{c}{L^{2}}\frac{\|\boldsymbol{\mu}\|^{2}}{\|\Sigma\|}\right),$
where $c$ is a constant which depends only on $\gamma$ and $q_{+}$.
Theorem 6.1 establishes the test error bound for (sG) in the strong signal regime, which was not studied in the existing literature. We see that benign overfitting occurs if $\|\boldsymbol{\mu}\|^{2}\gg\|\Sigma\|$ is satisfied in addition to the assumptions of the theorem.
For the weak signal regime, we can show
Theorem 6.2 (Weak Signal Regime).
Consider model (sG). Let $\delta\in(0,\tfrac{1}{4})$. Suppose $n\geq C$, $\alpha\leq\frac{1}{8{\rm tr}(\Sigma)}$, $\sigma\leq 0.1\gamma\alpha\sqrt{\frac{{\rm tr}(\Sigma)}{m}}$
${\rm tr}(\Sigma)\geq C\max\left\{n\|\boldsymbol{\mu}\|^{2},n\|\Sigma\|_{F},n^{\frac{3}{2}}\|\Sigma\|,n^{\frac{3}{2}}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|,n\frac{\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|^{2}}{\|\boldsymbol{\mu}\|^{2}}\right\},$
and further suppose one of the following:
  1. (i)
    $\|\boldsymbol{\mu}\|^{2}\geq C\max\left\{\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|,\sqrt{n}\|\Sigma\|_{F},n\|\Sigma\|\right\}$,
  2. (ii)
    ${\rm tr}(\Sigma)\geq C\max\left\{n^{\frac{3}{2}}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|,n^{\frac{3}{2}}\|\Sigma\|_{F},n^{2}\|\Sigma\|\right\}$
for a sufficiently large constant $C$ which depends only on $\delta,\gamma,q_{+}$. Then, we have with probably at least $1-3\delta$, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, converges in direction, and the convergent direction $\hat{W}$ attains
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(x;\hat{W})<0)\leq\exp\left(-\frac{c}{L^{2}}\frac{n\|\boldsymbol{\mu}\|^{4}}{\|\Sigma\|{\rm tr}(\Sigma)}\right),$
where $c$ is a constant which depends only on $\gamma$ and $q_{+}$.
Theorem 6.2 establishes the test error bound for (sG) in the weak signal regime, which corresponds to the prior works (Frei et al. (2022); Xu and Gu (2023)), which were limited to isotropic mixtures. We see that benign overfitting occurs if $n\|\boldsymbol{\mu}\|^{4}\gg\|\Sigma\|{\rm tr}(\Sigma)$.
As a corollary of these results, we provide the sufficient conditions of benign overfitting when $\Sigma=I_{p}$:
Corollary 6.3.
Consider model (sG) with $\Sigma=I_{p}$. Then, with high probability, the gradient descent iterate (2) converges in direction, and benign overfitting occurs with the convergent direction $\hat{W}$ if either one of the following holds:
  1. (i)
    $\|\boldsymbol{\mu}\|\gtrsim\sqrt{n}$ and $n^{\frac{5}{4}}\|\boldsymbol{\mu}\|^{\frac{3}{2}}\lesssim p\lesssim\min\{\tfrac{\|\boldsymbol{\mu}\|^{4}}{n},n\|\boldsymbol{\mu}\|^{2}\}$,
  2. (ii)
    $\|\boldsymbol{\mu}\|\gtrsim\sqrt{n}$ and $n\|\boldsymbol{\mu}\|^{2}\lesssim p\lesssim\frac{\|\boldsymbol{\mu}\|^{4}}{n}$,
  3. (iii)
    $\max\{n\|\boldsymbol{\mu}\|^{2},n^{3}\}\lesssim p\ll n\|\boldsymbol{\mu}\|^{4}$.
We note that the sufficient condition (iii) in the corollary is directly comparable to the results studied in the prior works (Frei et al. (2022); Xu and Gu (2023)), while conditions (i) and (ii) were not identified in the existing literature. Comparing the sufficient condition (iii) with that of Xu and Gu (2023), we do not require $\boldsymbol{z}$ to have a bounded log-Sobolev constant, and we have a better upper bound on $p$ (Xu and Gu (2023) requires $p\log(mp)\lesssim n\|\boldsymbol{\mu}\|^{4}$). We also do not require $m\gtrsim\log n$, which was necessary in their work to ensure $\min\{q_{+},q_{-}\}\geq 1/3$ after random initialization of $a_{j}$’s. Although we do not consider random initialization of $a_{j}$’s, our work on the weak signal regime can incorporate the random initialization of the form $a_{j}\sim\mathrm{Unif}\left\{\pm\tfrac{1}{\sqrt{m}}\right\}$ without assuming a lower bound on $m$. We note that our lower bound on $p$ is slightly stronger than theirs which requires $p\gtrsim\max\{n\|\boldsymbol{\mu}\|^{2},n^{2}\log n\}$.
The theorem statements here are a concise versions for the purpose of presentation. Detailed versions of these theorems and the proofs are given in Appendix C.
7 Proof Sketch
Proof sketch of Theorem 4.1.
The proof is done by reducing the argument to that of binary linear classification, i.e. classifiers of the form $\hat{y}=\mathrm{sgn}(\langle\boldsymbol{x},\boldsymbol{w}\rangle)$. In fact, if all the neurons are activated for all $(\boldsymbol{x}_{i},y_{i}),i=1,\ldots,n$ at step $t$, then we have
$y_{i}f(\boldsymbol{x}_{i};W^{(t)})=\langle y_{i}\boldsymbol{\tilde{x}}_{i},\boldsymbol{w}^{(t)}\rangle,$
where $\boldsymbol{w}^{(t)}=\mathrm{vec}(W^{(t)})$. Therefore, if all the neurons stay activated after $t\geq T$, gradient descent iterate (2) becomes that of linear classifier, and the directional convergence follows from Theorem 3, Soudry et al. (2018).
Proof sketch of Theorem 4.4 and 4.9.
To establish sufficient conditions for the neuron activation, observe that, for any $i,j$, we have
$a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t+1)}-\boldsymbol{w}_{j}^{(t)}\rangle$ (10)
$=$ $\frac{\alpha}{m}\Big{[}\zeta(\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t)}\rangle)\|\boldsymbol{x}_{i}\|^{2}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))+$
$\qquad+\sum_{k\neq i}\zeta(\langle\boldsymbol{x}_{k},\boldsymbol{w}_{j}^{(t)}\rangle)\langle y_{k}\boldsymbol{x}_{k},y_{i}\boldsymbol{x}_{i}\rangle\exp(-y_{k}f(\boldsymbol{x}_{k};W^{(t)}))\Big{]}.$
The only source of negativity in (10) are $\langle y_{i}\boldsymbol{x}_{i},y_{k}\boldsymbol{x}_{k}\rangle,i\neq k$. We note that event $E$ implies
$\big{|}\langle y_{i}\boldsymbol{x}_{i},y_{k}\boldsymbol{x}_{k}\rangle-\|\boldsymbol{\mu}\|^{2}\big{|}\leq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z}),\forall i\neq k.$ (11)
So, the only negativity can be controlled by $2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$. Therefore, under the condition $2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ is small enough (Case 1 or Assumption 4.5 in Case 2), the update (10) should stay positive. Then, it should be enough to make all the neurons activated at step $t=1$, which can be done by taking initialization $\sigma$ sufficiently smaller than the step size $\alpha$ (Assumption 4.2 and 4.6).
Proof sketch of Theorem 5.1.
To apply Lemma B.1, it is necessary to understand the behavior of $\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}$. Letting $X=(\boldsymbol{x}_{1}^{\top},\boldsymbol{x}_{2}^{\top},\ldots,\boldsymbol{x}_{n}^{\top})^{\top}\in R^{n\times p}$ and recalling $\boldsymbol{\tilde{x}}_{ij}=a_{j}\zeta(y_{i}a_{j})\boldsymbol{x}_{i}$, we can write
$\tilde{X}$ $=\begin{pmatrix}a_{1}B_{1}X&a_{2}B_{2}X&\cdots&a_{m}B_{m}X\end{pmatrix},$ (12)
where $B_{j}\in\mathbb{R}^{n\times n}$ is the diagonal matrix whose diagonal entry $(B_{j})_{kk}=\zeta(a_{j}y_{k})$. Then we have
$\tilde{X}\tilde{X}^{\top}=\sum_{j\in J}a_{j}^{2}B_{j}XX^{\top}B_{j}.$ (13)
Notice that $B_{j}$’s are the same among $j\in J_{+}$ and $j\in J_{-}$, respectively. So, we can write $B_{+}=B_{j},j\in J_{+}$ and $B_{-}=B_{j},j\in J_{-}$. With $a_{j}^{2}=\tfrac{1}{m}$, we have
$\tilde{X}\tilde{X}^{\top}=\frac{|J_{+}|}{m}B_{+}XX^{\top}B_{+}+\frac{|J_{-}|}{m}B_{-}XX^{\top}B_{-}.$ (14)
The analysis of $\boldsymbol{y}(\tilde{X}\tilde{X}^{\top})^{-1}$ will be done through $XX^{\top}=\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}\boldsymbol{y}^{\top}+\boldsymbol{y}(Z\boldsymbol{\mu})^{\top}+Z\boldsymbol{\mu}\boldsymbol{y}^{\top}+ZZ^{\top}\approx\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}\boldsymbol{y}^{\top}+RI_{n}$ under sufficient near orthogonality of $\boldsymbol{z}_{i}$’s to each other and to $\boldsymbol{\mu}$.
The rest of the argument is split into the strong signal regime and the weak signal regime. The latter is easier since we have $XX^{\top}\approx RI_{n}$, which further implies $\tilde{X}\tilde{X}^{\top}\approx R(q_{+}B_{+}^{2}+q_{-}B_{-}^{2})$, which is a diagonal matrix. The strong signal case is more difficult to analyze, but this can be done by a careful analysis of the eigenvectors of $\|\boldsymbol{\mu}\|^{2}(q_{+}B_{+}\boldsymbol{y}\boldsymbol{y}^{\top}B_{+}+q_{-}B_{-}\boldsymbol{y}\boldsymbol{y}^{\top}B_{-})+R(q_{+}B_{+}^{2}+q_{-}B_{-}^{2})$.
8 Discussion
In this paper, we have established sufficient conditions of directional convergence of gradient descent in leaky ReLU two-layer neural networks and, as an application, also derived sufficient conditions of benign overfitting. Here we discuss directions for future research other than the obvious direction of extending to deeper neural networks:
Our assumptions of Theorem 6.1 and 6.2 are stronger than what was imposed in prior works on binary linear classification (Wang and Thrampoulidis (2022); Cao et al. (2021); Hashimoto et al. (2025)). This is due to the asymmetric nature of leaky ReLU activation, which leads to complications that are not present in binary linear classification. First, gradient descent iterates are more difficult to analyze. So, we had to impose additional assumptions to obtain directional convergence. Second, due to the more complicated structure of $\tilde{X}\tilde{X}^{\top}$, we had to base our analysis of $(\tilde{X}\tilde{X}^{\top})^{-1}$ on cruder techniques than prior work on binary linear classification. Given prior work on gradient flow (Lyu and Li (2020); Ji and Telgarsky (2020)), we conjecture that directional convergence can be shown only assuming linear separability of the training data, which is the only assumption necessary for binary linear classification (Soudry et al. (2018)).
Our analysis revealed that the equivalence of $\boldsymbol{\hat{w}}$ and the minimum norm least square estimator fails in the strong signal regime if $q_{\gamma}<\gamma$. We conjecture that $\boldsymbol{\hat{w}}$ would be equivalent to the minimum norm least square only using positive-labeled samples (or negative-labeled samples) in this case.
In this study, we did not introduce noise to the label $y$ as was done in some prior works on binary classification (Chatterji and Long (2021); Wang and Thrampoulidis (2022); Frei et al. (2022); Xu and Gu (2023); Hashimoto et al. (2025)). Introducing noise to the label would likely not change the result for the weak signal regime since in this regime the signal $\boldsymbol{\mu}$ is negligible. However, we conjecture that the convergent direction would be very different in the strong signal regime with or without noise, as was revealed in Hashimoto et al. (2025) for binary linear classification, and hence require more involved argument.
Appendix A Proof of Directional Convergence
A.1 Proof of Theorem 4.1
Lemma A.1.
Consider (M). If all the neurons are activated at step $t$, i.e. $a_{j}\langle\boldsymbol{w}_{j}^{(t)},y_{i}\boldsymbol{x}_{i}\rangle>0$ holds for all $i\in I$ and $j\in J$, then
$\boldsymbol{w}_{j}^{(t+1)}-\boldsymbol{w}_{j}^{(t)}=\alpha\sum_{i\in I}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))a_{j}\zeta(a_{j}y_{i})y_{i}\boldsymbol{x}_{i}.$
Proof.
By (2), we have
$\boldsymbol{w}_{j}^{(t+1)}-\boldsymbol{w}_{j}^{(t)}$ $=-\alpha\sum_{i}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))(-y_{i})\nabla_{\boldsymbol{w}_{j}}\left\{\sum_{k}a_{k}\phi(\langle\boldsymbol{w}_{k},\boldsymbol{x}_{i}\rangle)\right\}_{W=W^{(t)}}$
$=\alpha\sum_{i}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))y_{i}a_{j}\nabla_{\boldsymbol{w}_{j}}\phi(\langle\boldsymbol{w}_{j},\boldsymbol{x}_{i}\rangle)_{\boldsymbol{w}_{j}=\boldsymbol{w}_{j}^{(t)}}$
$=\alpha\sum_{i}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))a_{j}\zeta(a_{j}y_{i})y_{i}\boldsymbol{x}_{i},$
where the last equality is due to the assumption $a_{j}y_{i}\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{x}_{i}\rangle>0$ which implies $\mbox{sgn}(\langle\boldsymbol{w}_{j}^{(t)},\boldsymbol{x}_{i}\rangle)=\mbox{sgn}(a_{j}y_{i})$. ∎
Remark.
Lemma A.1 implies that the gradient descent update is same within $j\in J_{+}$ and $j\in J_{-}$, respectively.
Lemma A.2.
Suppose $W=(\boldsymbol{w}_{1},\ldots,\boldsymbol{w}_{m})$ has $\boldsymbol{w}_{j}=\boldsymbol{w}_{+}$ if $j\in J_{+}$ and $\boldsymbol{w}_{j}=\boldsymbol{w}_{-}$ if $j\in J_{-}$. Then, $f(\boldsymbol{x};W)$ has the linear decision boundary defined by $\boldsymbol{\bar{w}}:=\frac{|J_{+}|}{\sqrt{m}}\boldsymbol{w}_{+}-\frac{|J_{-}|}{\sqrt{m}}\boldsymbol{w}_{-}$, i.e. $\mathrm{sign}(f(\boldsymbol{x};W))=\mathrm{sign}(\langle\boldsymbol{x},\boldsymbol{\bar{w}}\rangle)$.
Proof.
We consider several cases separately:
Case $\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle\geq 0$ and $\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle\geq 0$: we have
$f(\boldsymbol{x};\hat{W})=\frac{|J_{+}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle-\frac{|J_{-}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle=\langle\boldsymbol{x},\boldsymbol{\bar{w}}\rangle.$
Case $\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle\geq 0$ and $\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle<0$: we have
$f(\boldsymbol{x};\hat{W})=\frac{|J_{+}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle-\frac{\gamma|J_{-}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle\geq\frac{|J_{+}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle-\frac{|J_{-}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle=\langle\boldsymbol{x},\boldsymbol{\bar{w}}\rangle\geq 0.$
Case $\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle<0$ and $\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle\geq 0$: we have
$0>\langle\boldsymbol{x},\boldsymbol{\bar{w}}\rangle=\frac{|J_{+}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle-\frac{|J_{-}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle\geq\frac{\gamma|J_{+}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle-\frac{|J_{-}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle=f(\boldsymbol{x};\hat{W}).$
Case $\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle<0$ and $\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle<0$: we have
$f(\boldsymbol{x};\hat{W})=\gamma\left(\frac{|J_{+}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{+}\rangle-\frac{|J_{-}|}{\sqrt{m}}\langle\boldsymbol{x},\boldsymbol{w}_{-}\rangle\right)=\gamma\langle\boldsymbol{x},\boldsymbol{\bar{w}}\rangle.$
Therefore, we have confirmed $\mathrm{sign}(f(\boldsymbol{x}:\hat{W}))=\mathrm{sign}(\langle\boldsymbol{x},\boldsymbol{\bar{w}}\rangle)$ holds for all the cases above. ∎
Proof of Theorem 4.1.
By the assumption that all the neurons stay activated after $t\geq T$, the directional convergence follows from Corollary 8 of Soudry et al. [2018].
For completeness and to show the remaining claims, we provide the complete proof of the directional convergence here.
First note that $a_{j}\phi(\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t)})\rangle=a_{j}\zeta(a_{j}y_{i})\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t)}\rangle$ holds for all $i\in I,j\in J,t\geq T$ by the assumption that all the neurons are activated after $t\geq T$. Then, we have for $t\geq T$,
$y_{i}f(\boldsymbol{x}_{i};W^{(t)})$ $=\sum_{j}a_{j}y_{i}\phi(\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t)}\rangle)$
$=\sum_{j}a_{j}y_{i}\zeta(a_{j}y_{i})\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t)}\rangle$
$=\sum_{j}\langle a_{j}\zeta(a_{j}y_{i})y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t)}\rangle$
$=\sum_{j}\langle y_{i}\boldsymbol{\tilde{x}}_{ij},\boldsymbol{w}_{j}^{(t)}\rangle$
$=\langle y_{i}\boldsymbol{\tilde{x}}_{i},\boldsymbol{w}^{(t)}\rangle,$
where the fourth equality follows from the definition of $\boldsymbol{\tilde{x}}_{i}$ and $\boldsymbol{w}^{(t)}=\mathrm{vec}(W^{(t)})$.
With this, the gradient descent updates in Lemma A.1 can be rewritten as
$\boldsymbol{w}_{j}^{(t+1)}-\boldsymbol{w}_{j}^{(t)}=\alpha\sum_{i}\exp\left(-\langle y_{i}\boldsymbol{\tilde{x}}_{i},\boldsymbol{w}^{(t)}\rangle\right)a_{j}\zeta(a_{j}y_{i})y_{i}\boldsymbol{x}_{i}.$
Therefore, we have
$\boldsymbol{w}^{(t+1)}-\boldsymbol{w}^{(t)}=\alpha\sum_{i\in I}\exp(-\langle y_{i}\boldsymbol{\tilde{x}}_{i},\boldsymbol{w}^{(t)}\rangle)y_{i}\boldsymbol{\tilde{x}}_{i},$ (15)
which can be viewed as the gradient descent iterates of binary linear classifier of the dataset $\{(\boldsymbol{\tilde{x}}_{i},y_{i})\}_{i\in I}\subset\mathbb{R}^{mp}\times\{\pm 1\}$.
Therefore, gradient descent iterates $\boldsymbol{w}^{(t)}$ converges in direction if the dataset $\{(\boldsymbol{\tilde{x}}_{i},y_{i})\}_{i\in I}$ is linearly separable and $\alpha$ is sufficiently small due to Theorem 3 of Soudry et al. [2018]. We now provide additional details verifying that this result is applicable in our setting.
It is easy to see linear separability of $\{(\boldsymbol{\tilde{x}}_{i},y_{i})\}_{i\in I}$ by expanding the following inner product
$\langle y_{i}\boldsymbol{\tilde{x}}_{i},\boldsymbol{w}^{(t)}\rangle=\sum_{j\in J}\langle y_{i}a_{j}\zeta(a_{j}y_{i})\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t)}\rangle=\sum_{j\in J}\zeta(a_{j}y_{i})a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t)}\rangle.$ (16)
Now it is clear that the inner product is positive by the assumption that all the neurons are activated at time $t\geq T$.
In addition, $u\mapsto e^{-u}$ is 1-Lipschitz on $(0,\infty)$ and (16) implies that $\langle y_{i}\boldsymbol{\tilde{x}}_{i},\boldsymbol{w}^{(t)}\rangle$ stays in this region. Therefore, Theorem 3 of Soudry et al. [2018] can be used with $\beta=1$.
Since $\{(\boldsymbol{\tilde{x}}_{i},y_{i})\}_{i\in I}$ is linearly separable and satisfies $\alpha\sigma_{max}^{2}(\tilde{X})<2$, $\lim_{t\to\infty}\|\boldsymbol{w}^{(t)}\|=\infty$ follows from Lemma 1 of Soudry et al. [2018], $\mathcal{L}(W^{(t)})=O(t^{-1})$ follows from Theorem 5 of Soudry et al. [2018], and the directional convergence follows from Theorem 3 of Soudry et al. [2018], respectively.
Now we turn to the characterization of the convergent direction. First, we show that not only $\|\boldsymbol{w}^{(t)}\|$ but also $\|\boldsymbol{w}_{j}^{(t)}\|$ diverges to infinity for all $j\in J$. To see this, note that by Lemma A.1, we have for $t\geq T$
$\boldsymbol{w}_{j}^{(t)}-\boldsymbol{w}_{j}^{(T)}=\alpha\sum_{i}\left(\sum_{s=T}^{t-1}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(s)}))\right)\zeta(a_{j}y_{i})a_{j}y_{i}\boldsymbol{x}_{i}.$ (17)
Since $\|\boldsymbol{w}^{(t)}\|$ is diverging to infinity, we must have $\|\boldsymbol{w}_{j}^{(t)}\|$ diverging to infinity for some $j$, and hence $\sum_{s=T}^{t-1}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(s)}))\to\infty$ as $t\to\infty$ for some $i$. Denote such $i$ as $i_{*}$.
Again by Lemma A.1, we have for any $j$,
$\langle\boldsymbol{w}_{j}^{(t)}-\boldsymbol{w}_{j}^{(T)},\boldsymbol{w}_{j}^{(T)}\rangle$ $=\alpha\sum_{i}\left(\sum_{s=T}^{t-1}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(s)}))\right)\zeta(a_{j}y_{i})a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(T)}\rangle$ (18)
$\geq\alpha\sum_{s=T}^{t-1}\exp(-y_{i_{*}}f(\boldsymbol{x}_{i_{*}};W^{(s)}))\zeta(a_{j}y_{i_{*}})a_{j}\langle y_{i_{*}}\boldsymbol{x}_{i_{*}},\boldsymbol{w}_{j}^{(T)}\rangle\to\infty,$
where the inequality is due to the assumption that all the neurons are activated at $s\geq T$ and the divergence is due to $\sum_{s=T}^{t-1}\exp(-y_{i_{*}}f(\boldsymbol{x}_{i_{*}};W^{(s)}))\to\infty$ and the fact that all neurons are activated at time $T$, implying $\langle y_{i_{*}}\boldsymbol{x}_{i_{*}},\boldsymbol{w}_{j}^{(T)}\rangle\neq 0$. Since this holds for any $j$, we must have $\|\boldsymbol{w}_{j}\|$ diverging to infinity for all $j\in J$.
By Theorem 3 of Soudry et al. [2018], the convergent direction is equal to that of the solution to the following optimization problem:
$\text{Minimize}\quad\|\boldsymbol{w}\|^{2}\quad\text{s.t.}\quad\langle y_{i}\boldsymbol{\tilde{x}}_{i},\boldsymbol{w}\rangle\geq 1,\forall i\in I,$ (19)
which is equivalently written as
$\text{Minimize}\quad\sum_{j\in J}\|\boldsymbol{w}_{j}\|^{2}\quad\text{s.t.}\quad\sum_{j\in J}\langle a_{j}y_{i}\zeta(a_{j}y_{i})\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle\geq 1,\forall i\in I.$ (20)
Recalling that the gradient descent updates are same within $j\in J_{+}$ and $j\in J_{-}$, respectively, and that $\|\boldsymbol{w}_{j}^{(t)}\|$ diverges to infinity for all $j\in J$, the direction that $\boldsymbol{w}_{j}^{(t)}$ converges to must be the same for all $j\in J_{+}$ and the same for all $j\in J_{-}$. Thus the solution to the optimization problem (20) must have $\boldsymbol{w}_{j}=\boldsymbol{w}_{+}$ for all $j\in J_{+}$ and $\boldsymbol{w}_{j}=\boldsymbol{w}_{-}$ for all $j\in J_{-}$. With this, (20) can be further rewritten as the desired optimization problem (5).
Finally, the claim about the linear decision boundary follows from Lemma A.2
Before proceeding to the subsequent subsections where we prove Theorem 4.4 and 4.9, we introduce some ancillary results:
We first note that event $E$ implies (11) and the following inequality:
$(1-\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{min}^{2}(\boldsymbol{z})\}\leq\|y_{i}\boldsymbol{x}_{i}\|^{2}\leq(1+\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\},\forall i.$ (21)
We also note that the definition of $\rho$ implies
$\max_{i}|f(\boldsymbol{x}_{i};W^{(0)})|\leq\rho.$ (22)
By (11) and (21), we obtain the following lemma, which is going to be useful to establish sufficient conditions for $\alpha\sigma_{max}^{2}(\tilde{X})<2$:
Lemma A.3.
$\sigma_{max}^{2}(\tilde{X})\leq(1+\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}+(n-1)\left[\|\boldsymbol{\mu}\|^{2}+\{2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})\}\right]$
Proof.
Note that
$\left|\langle\boldsymbol{\tilde{x}_{i}},\boldsymbol{\tilde{x}_{k}}\rangle\right|$ $=\left|\sum_{j}a_{j}^{2}\zeta(a_{j}y_{i})\zeta(a_{j}y_{k})\langle\boldsymbol{x}_{i},\boldsymbol{x}_{k}\rangle\right|$ (23)
$\leq\sum_{j}\frac{\zeta(a_{j}y_{i})\zeta(a_{j}y_{k})}{m}\left|\langle\boldsymbol{x}_{i},\boldsymbol{x}_{k}\rangle\right|$
$\leq\left|\langle\boldsymbol{x}_{i},\boldsymbol{x}_{k}\rangle\right|,$
where the last inequality is due to $\gamma^{2}\leq\zeta(a_{j}y_{i})\zeta(a_{j}y_{k})\leq 1$.
With this, (11), and (21), for any $\boldsymbol{v}\in S^{n-1}$, we have
$\|\tilde{X}^{\top}\boldsymbol{v}\|^{2}$ $=\sum_{i}v_{i}^{2}\|\boldsymbol{\tilde{x}}_{i}\|^{2}+\sum_{i,k:i\neq k}v_{i}v_{k}\langle\boldsymbol{\tilde{x}}_{i},\boldsymbol{\tilde{x}}_{k}\rangle$
$\leq(1+\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}\sum_{i}v_{i}^{2}$
$\qquad+\left[\|\boldsymbol{\mu}\|^{2}+\{2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})\}\right]\sum_{i,k:i\neq k}|v_{i}v_{k}|$
$\leq(1+\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}$
$\qquad+(n-1)\left[\|\boldsymbol{\mu}\|^{2}+\{2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})\}\right],$
where in the last inequality we used $\sum_{i,k:i\neq k}|v_{i}v_{k}|\leq n-1$, which follows from the Cauchy-Schwartz inequality. ∎
A.2 Proof of Theorem 4.4
Lemma A.4.
Suppose event $E$ holds with $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ and Assumption 4.2. Then, all the neurons stay activated for $t\geq 1$.
Proof of Lemma A.4.
Under $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$, the gradient descent updates (10) stays positive. Therefore, it is enough to confirm all the neurons are activated at $t=1$.
Again by (10), (21), (22), $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$,
$a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(1)}\rangle$ (24)
$\geq$ $\frac{1}{m}\Big{[}\alpha\gamma(1-\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{min}^{2}(\boldsymbol{z})\}$
×$\qquad\qquad\qquad\qquad\times\exp\{-\sqrt{m}\sigma\max_{i}\|\boldsymbol{x}_{i}\|\}-\sqrt{m}\sigma\max_{i}\|\boldsymbol{x}_{i}\|\Big{]}$
$\geq$ $\frac{1}{m}\Big{[}\alpha\gamma(1-\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{min}^{2}(\boldsymbol{z})\}\exp(-\rho)-\rho\Big{]}>0.$
Proof of Theorem 4.4.
By Lemma A.4, all the neurons stay activated after $t\geq 1$.
By Assumption 4.3, it suffices to show $\sigma_{max}^{2}(\tilde{X})\leq 2(n\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z}))$ to ensure $\alpha\sigma_{max}^{2}(\tilde{X})<2$.
By Lemma A.3 and $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ we have
$\sigma_{max}^{2}(\tilde{X})$ $\leq(1+\theta_{2})(\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z}))+2(n-1)\|\boldsymbol{\mu}\|^{2}$ (25)
$\leq 2(\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z}))+2(n-1)\|\boldsymbol{\mu}\|^{2}$
$\leq 2(n\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})).$
Therefore, the directional convergence follows by Theorem 4.1. ∎
A.3 Proof of Theorem 4.9
First, it is useful to recognize the following inequality which follows from (10), (11), and (21):
$\frac{\alpha}{m}\Big{[}\gamma(1-\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{min}^{2}(\boldsymbol{z})\}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))$ (26)
$\quad+\sum_{k:k\neq i}\zeta(\langle\boldsymbol{x}_{k},\boldsymbol{w}_{j}^{(t)}\rangle)\{\|\boldsymbol{\mu}\|^{2}-2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})-\theta_{1}R_{max}^{2}(\boldsymbol{z})\}\exp(-y_{k}f(\boldsymbol{x}_{k};W^{(t)}))\Big{]}$
$\leq a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t+1)}-\boldsymbol{w}_{j}^{(t)}\rangle$
$\leq\frac{\alpha}{m}\Big{[}(1+\theta_{2})\{\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))$
$\quad\qquad+\sum_{k:k\neq i}\{\|\boldsymbol{\mu}\|^{2}+2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})\}\exp(-y_{k}f(\boldsymbol{x}_{k};W^{(t)}))\Big{]}.$
We start from establishing neuron activation at $t=1$.
Lemma A.5.
Suppose event $E$, Assumption 4.5 and 4.6 with $\varepsilon_{1}+\varepsilon_{2}<1$. Then, we have for any $i,j$,
$a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(1)}\rangle>0,$
and hence all the neurons are activated at $t=1$.
Proof.
By (22), (26), and the assumptions, we have
$a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(1)}\rangle$ $\geq\frac{1}{m}\Big{[}\alpha\Big{\{}\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})\exp(-\rho)$
$\qquad\qquad-n\{2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})\}\exp(\rho)\Big{\}}-\rho\Big{]}$
$\geq\frac{1}{m}\left[(1-\varepsilon_{1})\alpha\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})\exp(-\rho)-\rho\right]$
$\geq\frac{1}{m}(1-\varepsilon_{1}-\varepsilon_{2})\alpha\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})\exp(-\rho)>0.$
Thus, all the neurons are activated at $t=1$ if $\varepsilon_{1}+\varepsilon_{2}<1$. ∎
Lemma A.6.
Suppose event $E$, Assumption 4.5, 4.6, 4.7, and 4.8 with $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$. Then we have
$\max_{k\neq i}\frac{\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(1)}))}{\exp(-y_{k}f(\boldsymbol{x}_{k};W^{(1)}))}\leq\exp(1)\quad\text{and}\quad\alpha\sigma_{max}^{2}(\tilde{X})<2.$
Proof.
By (22), (26), and the assumptions, we have
$a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(1)}\rangle$ $\leq\frac{1}{m}\Big{[}\alpha\Big{\{}(1+\theta_{2})\{n\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}\exp(\rho)$
$\qquad\qquad+n\{2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})\}\exp(\rho)\Big{\}}+\rho\Big{]}$
$\leq\frac{1}{m}\Big{[}\alpha\Big{\{}(1+\theta_{2})(1+\varepsilon_{3})R_{max}^{2}(\boldsymbol{z})+\varepsilon_{1}R_{max}^{2}(\boldsymbol{z})\Big{\}}\exp(\rho)+\rho\Big{]}$
$\leq\frac{\alpha}{m}(1+\theta_{2})(1+\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3})R_{max}^{2}(\boldsymbol{z})\exp(\rho)$
$\leq\frac{4\alpha R_{max}^{2}(\boldsymbol{z})\exp(\rho)}{m}$
$\leq\frac{1}{m},$
where the second inequality is due to Assumption 4.5 and 4.7, the third is due to Assumption 4.6, the fourth is due to $\varepsilon_{1}+\varepsilon_{2}<1$, and the last inequality is due to Assumption 4.8.
Therefore, with Lemma A.5, we have $0<y_{i}f(\boldsymbol{x}_{i};W^{(1)})\leq 1$, and hence
$\max_{k\neq i}\frac{\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(1)}))}{\exp(-y_{k}f(\boldsymbol{x}_{k};W^{(1)}))}\leq\max_{k}\exp(y_{k}f(\boldsymbol{x}_{k};W^{(1)}))\leq\exp(1).$
We now show $\alpha\sigma_{max}^{2}(\tilde{X})<2$. By Lemma A.3, we have
$\alpha\sigma_{max}^{2}(\tilde{X})$ $\leq\alpha\left\{(1+\theta_{2})\{n\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}+\varepsilon_{1}(1+\theta_{2})R_{max}^{2}(\boldsymbol{z})\right\}$
$\leq\alpha(1+\theta_{2})(1+\varepsilon_{1}+\varepsilon_{3})R_{max}^{2}(\boldsymbol{z})$
$\leq 4\alpha R_{max}^{2}(\boldsymbol{z})<2,$
where the first inequality is due to Assumption 4.5, the second inequality is due to Assumption 4.7, and the last line is due to $\varepsilon_{1}+\varepsilon_{3}<1$ and Assumption 4.8.
Finally, we prove that all the neurons stay activated after $t\geq 1$.
Lemma A.7.
Suppose event $E$, Assumption 4.5, 4.6, 4.7, and 4.8 with $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$, $C_{w}\varepsilon_{i}<\frac{1}{2},i=1,3$, and $\theta_{2}\leq\frac{1}{2}$. Then we have for any $i,j$
$a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t)}\rangle>0$
and
$\max_{k\neq i}\frac{\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))}{\exp(-y_{k}f(\boldsymbol{x}_{k};W^{(t)}))}\leq C_{w}.$ (27)
Proof.
We prove both claims simultaneously by induction. By Lemma A.5 and A.6, both claims hold at step $t=1$. Suppose both claims hold for $t\geq 1$.
We first show the first claim also holds for $t+1$. By (26), Assumption 4.5, and the induction hypothesis, we have
$a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t+1)}-\boldsymbol{w}_{j}^{(t)}\rangle$ $\geq\frac{\alpha}{m}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))$ (28)
×$\quad\times\Big{\{}\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})-C_{w}\varepsilon_{1}\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})\Big{\}}$
$=\frac{1}{m}\gamma(1-C_{w}\varepsilon_{1})(1-\theta_{2})\alpha R_{min}^{2}(\boldsymbol{z})\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))$
$\geq\frac{1}{4m}\gamma\alpha R_{min}^{2}(\boldsymbol{z})\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))>0.$
where the last line is due to $C_{w}\varepsilon_{1}\leq\frac{1}{2}$ and $\theta_{2}\leq\frac{1}{2}$. Therefore, all the neurons are active at step $t+1$.
Next, we prove (27). By (26), the assumptions, and the induction hypothesis, we have
$a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t+1)}-\boldsymbol{w}_{j}^{(t)}\rangle$ (29)
$\leq$ $\frac{\alpha}{m}\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))$
×$\quad\times\Big{[}(1+\theta_{2})\big{(}1+\frac{\varepsilon_{3}}{n}\big{)}R_{max}^{2}(\boldsymbol{z})+C_{w}\varepsilon_{3}R_{max}^{2}(\boldsymbol{z})+C_{w}\varepsilon_{1}(1+\theta_{2})R_{max}^{2}(\boldsymbol{z})\Big{]}$
$\leq$ $\frac{1}{m}\Big{\{}1+\frac{\varepsilon_{3}}{n}+C_{w}\varepsilon_{3}+C_{w}\varepsilon_{1}\Big{\}}(1+\theta_{2})\alpha R_{max}^{2}(\boldsymbol{z})\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))$
$\leq$ $\frac{1}{m}6\alpha R_{max}^{2}(\boldsymbol{z})\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)})),$
where the first inequality is due to Assumption 4.5 and 4.7, the last is due to the assumptions on $\varepsilon_{i}$.
Since all the neurons are activated at both $t$ and $t+1$, we have
$a_{j}y_{i}\phi(\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t+1)}\rangle)-a_{j}y_{i}\phi(\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t)}\rangle)=\zeta(a_{j}y_{i})a_{j}\langle y_{i}\boldsymbol{x}_{i},\boldsymbol{w}_{j}^{(t+1)}-\boldsymbol{w}_{j}^{(t)}\rangle.$
With this, (28) and (29), we have
$\frac{\gamma^{2}}{4}\alpha R_{min}^{2}(\boldsymbol{z})\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))$ (30)
$\qquad\qquad\leq y_{i}f(\boldsymbol{x}_{i};W^{(t+1)})-y_{i}f(\boldsymbol{x}_{i};W^{(t)})$
$\qquad\qquad\qquad\qquad\leq 6\alpha R_{max}^{2}(\boldsymbol{z})\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)})).$
Now we fix $i\neq k$ and let $A_{t}=\frac{\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))}{\exp(-y_{k}f(\boldsymbol{x}_{k};W^{(t)}))}$. By induction hypothesis, we have $A_{t}\leq C_{w}$. By (30), we have
$A_{t+1}$ $\leq A_{t}\exp\Big{[}-\frac{\gamma^{2}}{4}\alpha R_{min}^{2}(\boldsymbol{z})\exp(-y_{i}f(\boldsymbol{x}_{i};W^{(t)}))$
$\qquad\qquad\qquad+6\alpha R_{max}^{2}(\boldsymbol{z})\exp(-y_{k}f(\boldsymbol{x}_{k};W^{(t)}))\Big{]}$ (31)
$=A_{t}\exp\Big{[}-\frac{\gamma^{2}}{4}\alpha R_{min}^{2}(\boldsymbol{z})\exp(-y_{k}f(\boldsymbol{x}_{k};W^{(t)}))\Big{\{}A_{t}-\frac{24R_{max}^{2}(\boldsymbol{z})}{\gamma^{2}R_{min}^{2}(\boldsymbol{z})}\Big{\}}\Big{]}.$ (32)
If $A_{t}\geq\frac{24R_{max}^{2}(\boldsymbol{z})}{\gamma^{2}R_{min}^{2}(\boldsymbol{z})}$, then $A_{t+1}\leq A_{t}$ immediately follows from (32).
If $A_{t}<\frac{24R_{max}^{2}(\boldsymbol{z})}{\gamma^{2}R_{min}^{2}(\boldsymbol{z})}$, then (31) and the fact that all neurons are activated at time $t$, yielding $y_{i}f(x_{i};W^{(t)})>0$ for all $i$, implies
$A_{t+1}$ $\leq\frac{24R_{max}^{2}(\boldsymbol{z})}{\gamma^{2}R_{min}^{2}(\boldsymbol{z})}\exp\Big{[}6\alpha R_{max}^{2}(\boldsymbol{z})\Big{]}$
$\leq\frac{24R_{max}^{2}(\boldsymbol{z})}{\gamma^{2}R_{min}^{2}(\boldsymbol{z})}\exp(1)\leq C_{w},$
where the last line holds from Assumption 4.8. This concludes the induction argument proving the desired claims. ∎
Proof of Theorem 4.9.
The conclusion now follows immediately from Lemma A.6, A.7, and Theorem 4.1. ∎
Appendix B Proof of Theorem 5.1
The proof of Theorem 5.1 is done by first establishing the equivalence of the maximum margin vector and the minimum norm least square estimator,i.e. $\boldsymbol{\hat{w}}=\tilde{X}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$, and carefully analyzing the behavior.
To prove $\boldsymbol{\hat{w}}=\tilde{X}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$, we rely on the "proliferation of support vector" phenomenon due to Lemma 1, Hsu et al. [2021].
Lemma B.1.
If $\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}>0$ for all $i$, then
$\boldsymbol{\hat{w}}=\tilde{X}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}.$
Proof.
By Lemma 1, Hsu et al. [2021], the assumption implies all $(\boldsymbol{\tilde{x}}_{i},y_{i}),i=1,2,\ldots,n$ are support vectors. Under this condition, the optimization problem defining $\boldsymbol{\hat{w}}$ is equivalent to
$\boldsymbol{\hat{w}}=\operatorname*{arg\,min}\|\boldsymbol{w}\|^{2},\quad\text{subject to $\langle\boldsymbol{w},y_{i}\boldsymbol{\tilde{x}}_{i}\rangle=1$},$
and is also equivalent to the optimization problem defining the least square estimator. ∎
As noted in Section 7, the analysis is done by establishing $\tilde{X}\tilde{X}^{\top}\approx\|\boldsymbol{\mu}\|^{2}(q_{+}B_{+}\boldsymbol{y}\boldsymbol{y}^{\top}B_{+}+q_{-}B_{-}\boldsymbol{y}\boldsymbol{y}^{\top}B_{-})+R(q_{+}B_{+}^{2}+q_{-}B_{-}^{2})$.
Lemma B.2.
If event $\tilde{E}_{1}\cap\tilde{E}_{2}$ holds, then the following inequality holds
$\left\|XX^{\top}-\left(\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}\boldsymbol{y}^{\top}+RI_{n}\right)\right\|\leq\tilde{\varepsilon}R,$
where $\tilde{\varepsilon}:=\max\{\tilde{\varepsilon}_{1},\sqrt{n}\tilde{\varepsilon}_{2}\}$.
Proof.
Noting event $\tilde{E}_{2}$ implies $\|\boldsymbol{y}(Z\boldsymbol{\mu})^{\top}\|\leq\frac{\tilde{\varepsilon}_{2}}{3}R$ and $XX^{\top}=\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}\boldsymbol{y}^{\top}+\boldsymbol{y}(Z\boldsymbol{\mu})^{\top}+(Z\boldsymbol{\mu})\boldsymbol{y}^{\top}+ZZ^{\top}$, we have
$\left\|XX^{\top}-\left(\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}\boldsymbol{y}^{\top}+RI_{n}\right)\right\|$ $\leq\|\boldsymbol{y}(Z\boldsymbol{\mu})^{\top}\|+\|(Z\boldsymbol{\mu})\boldsymbol{y}^{\top}\|+\|ZZ^{\top}-RI_{n}\|$
$\leq\tfrac{\tilde{\varepsilon}}{3}R+\tfrac{\tilde{\varepsilon}}{3}R+\tfrac{\tilde{\varepsilon}}{3}R=\tilde{\varepsilon}R.$
With (14), this establishes $\tilde{X}\tilde{X}^{\top}\approx\|\boldsymbol{\mu}\|^{2}(q_{+}B_{+}\boldsymbol{y}\boldsymbol{y}^{\top}B_{+}+q_{-}B_{-}\boldsymbol{y}\boldsymbol{y}^{\top}B_{-})+R(q_{+}B_{+}^{2}+q_{-}B_{-}^{2})$. It is convenient to introduce $B^{2}=q_{+}B_{+}^{2}+q_{-}B_{-}^{2}$ where $B\in\mathbb{R}^{n\times n}$ is a diagonal matrix with
$B_{ii}=\begin{cases}\sqrt{q_{+}+\gamma^{2}q_{-}}&\text{if $i\in I_{+}$}\\ \sqrt{q_{-}+\gamma^{2}q_{+}}&\text{if $i\in I_{-}$}.\end{cases}$
Lemma B.3.
If event $\tilde{E}_{1}\cap\tilde{E}_{2}$ holds, then we have
$\left\|B^{-1}\tilde{X}\tilde{X}^{\top}B^{-1}-A\right\|\leq\tilde{\varepsilon}q_{\gamma}^{-1}R$
where
$A:=\|\boldsymbol{\mu}\|^{2}B^{-1}(q_{+}B_{+}\boldsymbol{y}\boldsymbol{y}^{\top}B_{+}+q_{-}B_{-}\boldsymbol{y}\boldsymbol{y}^{\top}B_{-})B^{-1}+RI_{n}.$
To obtain approximation of $(\tilde{X}\tilde{X}^{\top})^{-1}$ from Lemma B.3, we need the following ancillary lemma:
Lemma B.4.
If $\|U-V\|\leq sL$ and $V\succeq LI_{n}$ for some $L>0$ and $s\in[0,\tfrac{1}{2}]$, then
$\|U^{-1}-V^{-1}\|\leq 2sL^{-1}.$
Proof.
Since $V\succeq LI_{n}$ for $L>0$, $V$ is invertible and $\|V^{-1}\|\leq L^{-1}$.
Then we have
$\|I_{n}-V^{-1}U\|\leq\|V^{-1}\|\|U-V\|\leq s.$
Letting $S=I_{n}-V^{-1}U$, we have $\|S\|\leq s\leq\tfrac{1}{2}$. So, $U$ is invertible and
$\|(V^{-1}U)^{-1}-I_{n}\|\leq\|S\|+\|S\|^{2}+\cdots\leq\frac{s}{1-s}\leq 2s.$
Combining with $\|V^{-1}\|\leq L^{-1}$, we conclude
$\|U^{-1}-V^{-1}\|\leq\|V^{-1}\|\|(V^{-1}U)^{-1}-I_{n}\|\leq 2sL^{-1}.$
Now we are ready to obtain the following approximation of $(\tilde{X}\tilde{X}^{\top})^{-1}$, which our remaining arguments heavily relies on:
Lemma B.5.
If event $\tilde{E}_{1}\cap\tilde{E}_{2}$ holds with $\tilde{\varepsilon}\leq\tfrac{q_{\gamma}}{2}$, then we have
$\|B(\tilde{X}\tilde{X}^{\top})^{-1}B-A^{-1}\|\leq\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}.$
Proof.
Note that the additional assumption implies $\tilde{\varepsilon}q_{\gamma}^{-1}\leq\tfrac{1}{2}$. Then, the desired conclusion follows from Lemma B.3 and B.4 by taking $U=B^{-1}\tilde{X}\tilde{X}^{\top}B^{-1},V=A,s=\tilde{\varepsilon}q_{\gamma}^{-1}$, and $L=R$. ∎
The next lemma establishes sufficient conditions of $\boldsymbol{\hat{w}}=\tilde{X}(\tilde{X}\tilde{X}^{\top})^{-1}y$. Before proceeding to the lemma, we note that the following inequality holds under $\tilde{E}_{3}$:
$(1-\tilde{\varepsilon}_{3}^{2})\frac{n^{2}}{4}\leq n_{+}n_{-}\leq\frac{n^{2}}{4}.$ (33)
Lemma B.6.
Suppose event $\bigcap_{i=1,2,3}\tilde{E}_{i}$ holds with $\tilde{\varepsilon}\leq\tfrac{q_{\gamma}}{2}$ and further assume one of the following conditions holds:
  1. (i)
    $q_{+}q_{-}=0$, $n\|\boldsymbol{\mu}\|^{2}<\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}}R$, and $\sqrt{n}\tilde{\varepsilon}\leq\frac{\gamma^{3}}{4}$,
  2. (ii)
    $q_{+}q_{-}\neq 0$, $n\|\boldsymbol{\mu}\|^{2}\leq\lambda R$, and $\sqrt{n}\tilde{\varepsilon}\leq\frac{q_{\gamma}^{2}}{\big{\{}(1+\tilde{\varepsilon}_{3})\lambda+2\big{\}}^{2}}$ for some $\lambda<\frac{q_{\gamma}}{\gamma(1-\gamma)(1+\tilde{\varepsilon}_{3})}$,
  3. (iii)
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq\lambda R$, and $\sqrt{n}\tilde{\varepsilon}\leq\frac{(1-\tilde{\varepsilon}_{3})q_{\gamma}^{2}}{4}\tilde{C}(\gamma,q_{+},\tilde{\varepsilon}_{3},\lambda)\frac{R}{n\|\boldsymbol{\mu}\|^{2}}$ for some $\lambda>0$, where
    $\tilde{C}(\gamma,q_{+},\lambda):=\frac{q_{\gamma}-\gamma}{\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}}.$
Then, we have
$\boldsymbol{\hat{w}}=\tilde{X}(\tilde{X}\tilde{X}^{\top})^{-1}y.$
Proof.
By Lemma B.1, it suffices to show $\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}>0$ for all $i$.
By Lemma B.5, we have
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}$ (34)
$=$ $(B^{-1}\boldsymbol{y})^{\top}B(\tilde{X}\tilde{X}^{\top})^{-1}BB^{-1}y_{i}\boldsymbol{e}_{i}$
$\geq$ $(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}y_{i}\boldsymbol{e}_{i}-\|B^{-1}\boldsymbol{y}\|B_{ii}^{-1}2\tilde{\varepsilon}q_{\gamma}^{-1}R^{-1}$
$\geq$ $B_{ii}^{-1}(B^{-1}\boldsymbol{y})^{\top}A^{-1}y_{i}\boldsymbol{e}_{i}-2B_{ii}^{-1}q_{\gamma}^{-1}\tilde{\varepsilon}R^{-1}\sqrt{\frac{n_{+}}{q_{+}+\gamma^{2}q_{-}}+\frac{n_{-}}{q_{-}+\gamma^{2}q_{+}}}$
$\geq$ $B_{ii}^{-1}\Big{\{}(B^{-1}\boldsymbol{y})^{\top}A^{-1}y_{i}\boldsymbol{e}_{i}-2q_{\gamma}^{-3/2}\tilde{\varepsilon}R^{-1}\sqrt{n}\Big{\}},$
where the last inequality follows from the definition of $q_{\gamma}$ and $n_{+}+n_{-}=n$.
Let $A=\sum_{i=1}^{n}s_{i}\boldsymbol{u}_{i}\boldsymbol{u}_{i}^{\top}$ be the singular value decomposition of $A$ with $s_{1}\geq s_{2}\geq\ldots\geq s_{n}$. We split in cases.
We first consider the case either $q_{+}$ or $q_{-}=0$, i.e. either $|J_{+}|$ or $|J_{-}|=0$. Then we have $B=B_{+}$ if $q_{-}=0$ and $B=B_{-}$ if $q_{+}=0$, and hence $A=\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}\boldsymbol{y}^{\top}+RI_{n}$ either way. Then, we have $s_{1}=n\|\boldsymbol{\mu}\|^{2}+R$, $s_{i}=R,i\geq 2$, and can take $\boldsymbol{u}_{1}=\tfrac{\boldsymbol{y}}{\sqrt{n}}$. Thus,
$A^{-1}$ $=[n\|\boldsymbol{\mu}\|^{2}+R]^{-1}\tfrac{\boldsymbol{y}}{\sqrt{n}}\tfrac{\boldsymbol{y}^{\top}}{\sqrt{n}}+R^{-1}\sum_{i\geq 2}\boldsymbol{u}_{i}\boldsymbol{u}_{i}^{\top}$ (35)
$=R^{-1}\sum_{i=1}^{n}\boldsymbol{u}_{i}\boldsymbol{u}_{i}^{\top}-\frac{n\|\boldsymbol{\mu}\|^{2}}{[n\|\boldsymbol{\mu}\|^{2}+R]R}\frac{\boldsymbol{y}}{\sqrt{n}}\frac{\boldsymbol{y}^{\top}}{\sqrt{n}}$
$=R^{-1}\left(I_{n}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\boldsymbol{y}}{\sqrt{n}}\frac{\boldsymbol{y}^{\top}}{\sqrt{n}}\right).$
Let $y=\boldsymbol{1}_{+}-\boldsymbol{1}_{-}$, where $(\boldsymbol{1}_{+})_{i}=1$ if $i\in I_{+}$ and $(\boldsymbol{1}_{+})_{i}=0$ if $i\in I_{-}$ and $(\boldsymbol{1}_{-})_{i}=1$ if $i\in I_{-}$ and $(\boldsymbol{1}_{-})_{i}=0$ if $i\in I_{+}$.
Case $q_{+}=1,q_{-}=0$: In this case, we have $B=B_{+}$ as noted earlier. Since $\boldsymbol{y}=\boldsymbol{1}_{+}-\boldsymbol{1}_{-}$ and $B_{+}^{-1}\boldsymbol{y}=\boldsymbol{1}_{+}-\gamma^{-1}\boldsymbol{1}_{-}$, we have that
$(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}$ (36)
$=$ $R^{-1}\left\{B_{+}^{-1}\boldsymbol{y}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\langle B_{+}^{-1}\boldsymbol{y},\boldsymbol{y}\rangle}{n}\boldsymbol{y}\right\}^{\top}$
$=$ $R^{-1}\left\{\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)\boldsymbol{1}_{+}\right.$
$\qquad\qquad\left.-\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)\boldsymbol{1}_{-}\right\}^{\top}.$
Thus, we have
$(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}B_{+}^{-1}y_{i}\boldsymbol{e}_{i}$ (37)
$=$ $\begin{cases}R^{-1}\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right),&\text{if $i\in I_{+}$},\\ \gamma^{-1}R^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right),&\text{if $i\in I_{-}$}.\end{cases}$
Therefore, by the definition of $\tilde{E}_{3}$, (34) and (37), we have
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}$ (38)
$\geq$ $\begin{cases}R^{-1}\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\gamma^{-1})(1+\tilde{\varepsilon}_{3})}{2}-2\gamma^{-3}\sqrt{n}\tilde{\varepsilon}\right),&\text{if $i\in I_{+}$},\\ \gamma^{-1}R^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\gamma^{-1})(1+\tilde{\varepsilon}_{3})}{2}-2\gamma^{-3}\sqrt{n}\tilde{\varepsilon}\right),&\text{if $i\in I_{-}$}.\end{cases}$
Case $q_{+}=0,q_{-}=1$: In this case, we have $B=B_{-}$ as noted earlier. Since $\boldsymbol{y}=\boldsymbol{1}_{+}-\boldsymbol{1}_{-}$ and $B_{-}^{-1}\boldsymbol{y}=\gamma^{-1}\boldsymbol{1}_{+}-\boldsymbol{1}_{-}$, we have that
$(B_{-}^{-1}\boldsymbol{y})^{\top}A^{-1}$ (39)
$=$ $R^{-1}\left\{B_{-}^{-1}\boldsymbol{y}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\langle B_{-}^{-1}\boldsymbol{y},\boldsymbol{y}\rangle}{n}\boldsymbol{y}\right\}^{\top}$
$=$ $R^{-1}\left\{\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\gamma^{-1}n_{+}+n_{-}}{n}\right)\boldsymbol{1}_{+}\right.$
$\qquad\qquad\left.-\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\gamma^{-1}n_{+}+n_{-}}{n}\right)\boldsymbol{1}_{-}\right\}^{\top}.$
Thus, we have
$(B_{-}^{-1}\boldsymbol{y})^{\top}A^{-1}B_{-}^{-1}y_{i}\boldsymbol{e}_{i}$ (40)
$=$ $\begin{cases}\gamma^{-1}R^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\gamma^{-1}n_{+}+n_{-}}{n}\right),&\text{if $i\in I_{+}$}\\ R^{-1}\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{\gamma^{-1}n_{+}+n_{-}}{n}\right),&\text{if $i\in I_{-}$}.\end{cases}$
Therefore, the definition of $\tilde{E}_{3}$, (34) and (40), we have
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}$ (41)
$\geq$ $\begin{cases}\gamma^{-1}R^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\gamma^{-1})(1+\tilde{\varepsilon}_{3})}{2}-2\gamma^{-3}\sqrt{n}\tilde{\varepsilon}\right),&\text{if $i\in I_{+}$},\\ R^{-1}\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\gamma^{-1})(1+\tilde{\varepsilon}_{3})}{2}-2\gamma^{-3}\sqrt{n}\tilde{\varepsilon}\right),&\text{if $i\in I_{-}$}.\end{cases}$
Therefore, for both the cases above, it suffices to have
$\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{(1+\gamma^{-1})(1+\tilde{\varepsilon}_{3})}{2}<\frac{1}{2},\quad\text{and}\quad 2\gamma^{-3}\sqrt{n}\tilde{\varepsilon}\leq\frac{1}{2},$
which follows from the condition (i).
Case $q_{+}q_{-}\neq 0$: In this case, we have the span of $\{B^{-1}B_{+}\boldsymbol{y},B^{-1}B_{-}\boldsymbol{y}\}$ equal to the span of $\{\boldsymbol{u}_{1},\boldsymbol{u}_{2}\}$, and $s_{i}=R$ for $i\geq 3$.
To determine $\sum_{i=1,2}s_{i}^{-1}\boldsymbol{u}_{i}\boldsymbol{u}_{i}^{\top}$, it is convenient to use isomorphism between the span of $\{B^{-1}B_{+}\boldsymbol{y},B^{-1}B_{-}\boldsymbol{y}\}$ and $\mathbb{R}^{2}$ by $\boldsymbol{1}_{+}\leftrightarrow\sqrt{n_{+}}\boldsymbol{e}_{1}$ and $\boldsymbol{1}_{-}\leftrightarrow\sqrt{n_{-}}\boldsymbol{e}_{2}$. Under this isomorphism, we have
$B^{-1}B_{+}\boldsymbol{y}=\frac{1}{\sqrt{q_{+}+\gamma^{2}q_{-}}}\boldsymbol{1}_{+}-\frac{\gamma}{\sqrt{q_{-}+\gamma^{2}q_{+}}}\boldsymbol{1}_{-}\leftrightarrow\frac{\sqrt{n_{+}}}{\sqrt{q_{+}+\gamma^{2}q_{-}}}\boldsymbol{e}_{1}-\frac{\gamma\sqrt{n_{-}}}{\sqrt{q_{-}+\gamma^{2}q_{+}}}\boldsymbol{e}_{2},$ (42)
$B^{-1}B_{-}\boldsymbol{y}=\frac{\gamma}{\sqrt{q_{+}+\gamma^{2}q_{-}}}\boldsymbol{1}_{+}-\frac{1}{\sqrt{q_{-}+\gamma^{2}q_{+}}}\boldsymbol{1}_{-}\leftrightarrow\frac{\gamma\sqrt{n_{+}}}{\sqrt{q_{+}+\gamma^{2}q_{-}}}\boldsymbol{e}_{1}-\frac{\sqrt{n_{-}}}{\sqrt{q_{-}+\gamma^{2}q_{+}}}\boldsymbol{e}_{2}.$
Thus, $\sum_{i=1,2}s_{i}\boldsymbol{u}_{i}\boldsymbol{u}_{i}^{\top}$ is regarded as a $2\times 2$ matrix
$\begin{pmatrix}n_{+}\|\boldsymbol{\mu}\|^{2}+R&-\dfrac{\gamma\sqrt{n_{+}n_{-}}\|\boldsymbol{\mu}\|^{2}}{\sqrt{(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})}}\\ -\dfrac{\gamma\sqrt{n_{+}n_{-}}\|\boldsymbol{\mu}\|^{2}}{\sqrt{(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})}}&n_{-}\|\boldsymbol{\mu}\|^{2}+R\end{pmatrix}.$ (43)
Thus, $\sum_{i=1,2}s_{i}^{-1}\boldsymbol{u}_{i}\boldsymbol{u}_{i}^{\top}$ is regarded as
$\begin{pmatrix}n_{+}\|\boldsymbol{\mu}\|^{2}+R&-\dfrac{\gamma\sqrt{n_{+}n_{-}}\|\boldsymbol{\mu}\|^{2}}{\sqrt{(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})}}\\ -\dfrac{\gamma\sqrt{n_{+}n_{-}}\|\boldsymbol{\mu}\|^{2}}{\sqrt{(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})}}&n_{-}\|\boldsymbol{\mu}\|^{2}+R\end{pmatrix}^{-1}$ (44)
$=$ $d^{-1}\begin{pmatrix}n_{-}\|\boldsymbol{\mu}\|^{2}+R&\dfrac{\gamma\sqrt{n_{+}n_{-}}\|\boldsymbol{\mu}\|^{2}}{\sqrt{(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})}}\\ \dfrac{\gamma\sqrt{n_{+}n_{-}}\|\boldsymbol{\mu}\|^{2}}{\sqrt{(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})}}&n_{+}\|\boldsymbol{\mu}\|^{2}+R\end{pmatrix},$
where
$d=\{n_{+}\|\boldsymbol{\mu}\|^{2}+R\}\{n_{-}\|\boldsymbol{\mu}\|^{2}+R\}-\frac{\gamma^{2}n_{+}n_{-}\|\boldsymbol{\mu}\|^{4}}{(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})}.$
Since $B^{-1}\boldsymbol{y}\leftrightarrow\frac{\sqrt{n_{+}}}{\sqrt{q_{+}+\gamma^{2}q_{-}}}\boldsymbol{e}_{1}-\frac{\sqrt{n_{-}}}{\sqrt{q_{-}+\gamma^{2}q_{+}}}\boldsymbol{e}_{2}$ and $B^{-1}\boldsymbol{y}\perp\boldsymbol{u}_{i}$ for all $i\geq 3$, we have
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}$ (45)
$=$ $d^{-1}\left[\left\{\frac{n_{-}\|\boldsymbol{\mu}\|^{2}+R}{\sqrt{q_{+}+\gamma^{2}q_{-}}}-\frac{\gamma n_{-}\|\boldsymbol{\mu}\|^{2}}{\sqrt{q_{+}+\gamma^{2}q_{-}}(q_{-}+\gamma^{2}q_{+})}\right\}\boldsymbol{1}_{+}\right.$
$\quad\qquad\left.-\left\{\frac{n_{+}\|\boldsymbol{\mu}\|^{2}+R}{\sqrt{q_{-}+\gamma^{2}q_{+}}}-\frac{\gamma n_{+}\|\boldsymbol{\mu}\|^{2}}{\sqrt{q_{-}+\gamma^{2}q_{+}}(q_{+}+\gamma^{2}q_{-})}\right\}\boldsymbol{1}_{-}\right]^{\top}$
$=$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}$
×$\quad\times\Big{[}\sqrt{q_{+}+\gamma^{2}q_{-}}\big{\{}(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\big{\}}\boldsymbol{1}_{+}$
$\qquad-\sqrt{q_{-}+\gamma^{2}q_{+}}\big{\{}(q_{+}+\gamma^{2}q_{-}-\gamma)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\big{\}}\boldsymbol{1}_{-}\Big{]}^{\top}.$
Therefore, by (34) and (45) , we have for $i\in I_{+}$
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}$ (46)
$\geq$ $B_{ii}^{-1}\Big{[}\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}$
×$\qquad\times\Big{\{}\sqrt{q_{+}+\gamma^{2}q_{-}}\big{\{}(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\big{\}}$
$\qquad\qquad\quad-2q_{\gamma}^{-3/2}R^{-1}\sqrt{n}\tilde{\varepsilon}\Big{]},$
and for $i\in I_{-}$
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}$ (47)
$\geq$ $B_{ii}^{-1}\Big{[}\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}$
×$\qquad\times\Big{\{}\sqrt{q_{-}+\gamma^{2}q_{+}}\big{\{}(q_{+}+\gamma^{2}q_{-}-\gamma)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\big{\}}$
$\qquad\qquad\quad-2q_{\gamma}^{-3/2}R^{-1}\sqrt{n}\tilde{\varepsilon}\Big{]}.$
Under regime (ii), for $i\in I_{+}$ we have
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})$ $\leq q_{\gamma}(n_{+}+\|\boldsymbol{\mu}\|^{2})(n_{-}+\|\boldsymbol{\mu}\|^{2})$ (48)
$\leq q_{\gamma}\left(\frac{1+\tilde{\varepsilon}_{3}}{2}n\|\boldsymbol{\mu}\|^{2}+R\right)^{2}$
$\leq q_{\gamma}\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}R^{2}.$
We also have
$(q_{-}+\gamma^{2}q_{+}-\gamma)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R$ (49)
$\geq$ $q_{\gamma}R-\frac{\gamma(1-\gamma)(1+\tilde{\varepsilon}_{3})}{2}n\|\boldsymbol{\mu}\|^{2}$
$>$ $\frac{q_{\gamma}}{2}R.$
By (46), (48), and (49), we have
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}>B_{ii}^{-1}R^{-1}\left[\frac{q_{\gamma}^{\frac{1}{2}}}{2\big{\{}\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\big{\}}^{2}}-2q_{\gamma}^{-3/2}\sqrt{n}\tilde{\varepsilon}\right]\geq 0.$ (50)
The same conclusion analogously holds for $i\in I_{-}$ as well.
Under regime (iii), we have that
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})$ (51)
$=$ $(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\{n_{+}\|\boldsymbol{\mu}\|^{2}+R\}\{n_{-}\|\boldsymbol{\mu}\|^{2}+R\}-\gamma^{2}n_{+}n_{-}\|\boldsymbol{\mu}\|^{4}$
$=$ $(1-\gamma^{2})^{2}q_{+}q_{-}n_{+}n_{-}\|\boldsymbol{\mu}\|^{4}+(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\{n\|\boldsymbol{\mu}\|^{2}+R\}R$
$\leq$ $\Big{[}\frac{(1-\gamma^{2})^{2}}{16}n^{2}\|\boldsymbol{\mu}\|^{4}+q_{\gamma}\{n\|\boldsymbol{\mu}\|^{2}+R\}R\Big{]}$
$\leq$ $\Big{[}\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\Big{]}(n\|\boldsymbol{\mu}\|^{2})^{2}.$
With this and (46), we have for $i\in I_{+}$
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y_{i}\boldsymbol{e}_{i}$ (52)
$>$ $B_{ii}^{-1}\left[\frac{q_{\gamma}^{\frac{1}{2}}\{(q_{\gamma}-\gamma)(1-\tilde{\varepsilon}_{3})\frac{n}{2}\|\boldsymbol{\mu}\|^{2}\}}{\Big{\{}\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\Big{\}}(n\|\boldsymbol{\mu}\|^{2})^{2}}-2q_{\gamma}^{-\frac{3}{2}}\sqrt{n}\tilde{\varepsilon}R^{-1}\right]$
$=$ $(B_{ii}n\|\boldsymbol{\mu}\|^{2})^{-1}q_{\gamma}^{-\frac{3}{2}}\left[\frac{q_{\gamma}^{2}(q_{\gamma}-\gamma)(1-\tilde{\varepsilon}_{3})}{2\Big{\{}\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\Big{\}}}-2\sqrt{n}\tilde{\varepsilon}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\right]\geq 0.$
Similarly, the same lower bound holds for $i\in I_{-}$ from (47).
Remark.
[…]
With $\boldsymbol{\hat{w}}=\tilde{X}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}y$, by defining
$\boldsymbol{w}_{+}$ $=\frac{1}{\sqrt{m}}X^{\top}B_{+}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$ (53)
$=\frac{1}{\sqrt{m}}X^{\top}B_{+}\left(q_{+}B_{+}XX^{\top}B_{+}+q_{-}B_{-}XX^{\top}B_{-}\right)^{-1}\boldsymbol{y},$
$\boldsymbol{w}_{-}$ $=-\frac{1}{\sqrt{m}}X^{\top}B_{-}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$
$=-\frac{1}{\sqrt{m}}X^{\top}B_{-}\left(q_{+}B_{+}XX^{\top}B_{+}+q_{-}B_{-}XX^{\top}B_{-}\right)^{-1}\boldsymbol{y},$
we have $\boldsymbol{\hat{w}}_{j}=\boldsymbol{w}_{+}$ if $j\in J_{+}$ and $\boldsymbol{\hat{w}}_{j}=\boldsymbol{w}_{-}$ if $j\in J_{-}$.
By Lemma A.2, the network $f(x;\hat{W})$ has the linear decision boundary defined by
$\boldsymbol{\bar{w}}=\tfrac{|J_{+}|}{\sqrt{m}}\boldsymbol{w}_{+}-\tfrac{|J_{-}|}{\sqrt{m}}\boldsymbol{w}_{-}.$ (54)
By (53) and (54), we have
$\boldsymbol{\bar{w}}$ $=\left(q_{+}X^{\top}B_{+}+q_{-}X^{\top}B_{-}\right)(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$
This implies
$\|\boldsymbol{\bar{w}}\|^{2}$ $=\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\left(\sum_{j\in\{+,-\}}q_{j}B_{j}X\right)\left(\sum_{j\in\{+,-\}}q_{j}B_{j}X\right)^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y},$ (55)
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$ $=\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\left(q_{+}B_{+}X+q_{-}B_{-}X\right)\boldsymbol{\mu}$
$=\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)(\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}+Z\boldsymbol{\mu}).$ (56)
We need to upper bound $\|\boldsymbol{\bar{w}}\|^{2}$ and lower bound $\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$, respectively, in order to later establish a classification error bound of $\boldsymbol{\bar{w}}$.
To obtain an upper bound for $\|\boldsymbol{\bar{w}}\|^{2}$, the following inequality is useful.
Lemma B.7.
$\|\boldsymbol{\bar{w}}\|^{2}\leq 2\max\{q_{+},q_{-}\}\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}.$
Proof.
Since $(U+V)(U+V)^{\top}\succeq O$ and $(U-V)(U-V)^{\top}\succeq O$, we have
$2(UU^{\top}+VV^{\top})=(U+V)(U+V)^{\top}+(U-V)(U-V)^{\top}\succeq(U+V)(U+V)^{\top}.$
Letting $U=q_{+}B_{+}X$ and $V=q_{-}B_{-}X$, we have by (55)
$\|\boldsymbol{\bar{w}}\|^{2}$ $\leq 2\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}(q_{+}^{2}B_{+}XX^{\top}B_{+}+q_{-}^{2}B_{-}XX^{\top}B_{-})(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$
$\leq 2\max\{q_{+},q_{-}\}\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}(q_{+}B_{+}XX^{\top}B_{+}+q_{-}B_{-}XX^{\top}B_{-})(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$
$=2\max\{q_{+},q_{-}\}\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}.$
Lemma B.8.
Suppose all the assumptions of Lemma B.6 hold with $\tilde{\varepsilon}_{3}\leq\tfrac{1}{2}$ and $\lambda\geq\frac{\gamma}{4}$ in (iii). Then, there is a constant $\tilde{c}_{1}$ which only depends on $\gamma$ and $q_{+}$ such that we have
$\|\boldsymbol{\bar{w}}\|^{2}\leq\tilde{c}_{1}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}.$
Proof.
By Lemma B.7, it is reduced to obtaining an upper bound of $\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$. By Lemma B.5 and the definition of $q_{\gamma}$, we have
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$ $\leq(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\boldsymbol{y}+\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\left(\frac{n_{+}}{q_{+}+\gamma^{2}q_{-}}+\frac{n_{-}}{q_{-}+\gamma^{2}q_{+}}\right)$ (57)
$\leq(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\boldsymbol{y}+\frac{2q_{\gamma}^{-2}n\tilde{\varepsilon}}{R}.$
We split in cases.
Case $q_{+}=1,q_{-}=0$: By (36), we have
$(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}(B_{+}^{-1}\boldsymbol{y})$ (58)
$=$ $R^{-1}\left\{\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)n_{+}\right.$
$\qquad\qquad\left.+\gamma^{-1}\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)n_{-}\right\}.$
By (57), we have under the regime (i)
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}\;\leq$ $\;R^{-1}\left\{\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)n_{+}\right.$ (59)
[…]
$\leq R^{-1}\left\{(n_{+}+\gamma^{-2}n_{-})+2nq_{\gamma}^{-2}\tilde{\varepsilon}\right\}$
$\leq\left(\gamma^{-2}+2q_{\gamma}^{-2}\tilde{\varepsilon}\right)\frac{n}{R}$
$\leq\frac{3}{2\gamma^{2}}\frac{n}{R},$
where the last inequality is due to the condition on $\tilde{\varepsilon}$ of the regime (i).
Case $q_{+}=0,q_{-}=1$: Similarly, by (39) and (57), we obtain the same bound as the case $q_{+}=1,q_{-}=0$.
Case $q_{+}q_{-}\neq 0$: By (45), we have
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\boldsymbol{y}$ (60)
$=$ $\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}$
×[$\quad\times\big{[}\left\{\left(q_{-}+\gamma^{2}q_{+}-\gamma\right)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\right\}n_{+}$
+$\quad+\left\{\left(q_{+}+\gamma^{2}q_{-}-\gamma\right)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\right\}n_{-}\big{]}$
$\leq\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\big{\{}(1-\gamma)^{2}n_{+}n_{-}\|\boldsymbol{\mu}\|^{2}+nR\big{\}}$
$\leq\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}\left\{(1-\gamma)^{2}\frac{n^{2}}{4}\|\boldsymbol{\mu}\|^{2}+nR\right\},$
where the last line follows from (33).
Under the regime (ii), we first note
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})$
$=$ $(1-\gamma^{2})^{2}q_{+}q_{-}n_{+}n_{-}\|\boldsymbol{\mu}\|^{4}+(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\{n\|\boldsymbol{\mu}\|^{2}+R\}R$
$\geq$ $(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})R^{2}$
$\geq$ $q_{\gamma}^{2}R^{2}.$
With this and (60), we have
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\boldsymbol{y}$ $\leq(q_{\gamma}^{2}R^{2})^{-1}\left\{(1-\gamma)^{2}\frac{n^{2}}{4}\|\boldsymbol{\mu}\|^{2}+nR\right\}$ (61)
$\leq q_{\gamma}^{-2}\left\{1+\frac{(1-\gamma)^{2}}{4}\lambda\right\}\frac{n}{R}.$
Therefore, by (57), we have
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$ $\leq q_{\gamma}^{-2}\left\{1+\frac{(1-\gamma)^{2}}{4}\lambda+2\tilde{\varepsilon}\right\}\frac{n}{R}$ (62)
$\leq q_{\gamma}^{-2}\left\{2+\frac{(1-\gamma)^{2}}{4}\lambda\right\}\frac{n}{R}$
$\leq q_{\gamma}^{-2}\left\{2+\frac{1-\gamma}{4\gamma}q_{\gamma}\right\}\frac{n}{R}$
$\leq\frac{3}{\gamma q_{\gamma}^{2}}\frac{n}{R},$
where the second inequality is due to the condition on $\tilde{\varepsilon}$ of the regime (ii) and the third inequality is due to the upper bound on $\lambda$ of the regime (ii).
Under the regime (iii), we have
$d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\geq(1-\gamma^{2})^{2}q_{+}q_{-}n^{2}\|\boldsymbol{\mu}\|^{4}\geq(1-\gamma^{2})^{2}q_{+}q_{-}(1-\tilde{\varepsilon}_{3}^{2})\frac{n^{2}\|\boldsymbol{\mu}\|^{4}}{4},$
where the last inequality is due to (33).
With this and (60), we have
$(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\boldsymbol{y}$ (63)
$\leq$ $\left\{(1-\gamma^{2})q_{+}q_{-}(1-\tilde{\varepsilon}_{3}^{2})\frac{n^{2}\|\boldsymbol{\mu}\|^{4}}{4}\right\}^{-1}\left\{(1-\gamma)^{2}\frac{n^{2}}{4}\|\boldsymbol{\mu}\|^{2}+nR\right\}$
$\leq$ $\frac{(1-\gamma)^{2}+4\lambda^{-1}}{(1-\gamma^{2})^{2}q_{+}q_{-}(1-\tilde{\varepsilon}_{3}^{2})}\|\boldsymbol{\mu}\|^{-2}.$
Therefore, by (57), $\tilde{\varepsilon}_{3}\leq\tfrac{1}{2}$, and $\lambda\geq\tfrac{\gamma}{4}$, we have
$\boldsymbol{y}^{\top}(\tilde{X}\tilde{X}^{\top})^{-1}\boldsymbol{y}$ (64)
$\leq$ $\frac{(1-\gamma)^{2}+4\lambda^{-1}}{(1-\gamma^{2})^{2}q_{+}q_{-}(1-\tilde{\varepsilon}_{3}^{2})}\|\boldsymbol{\mu}\|^{-2}+\frac{2q_{\gamma}^{-2}n\tilde{\varepsilon}}{R}$
$\leq$ $\left[\frac{(1-\gamma)^{2}+4\lambda^{-1}}{(1-\gamma^{2})^{2}q_{+}q_{-}(1-\tilde{\varepsilon}_{3}^{2})}+\frac{(q_{\gamma}-\gamma)(1-\tilde{\varepsilon}_{3})}{4\big{\{}\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\big{\}}}\frac{2}{\sqrt{n}}\right]\|\boldsymbol{\mu}\|^{-2}$
$\leq$ $\left[\frac{4(1-\gamma)^{2}+64\gamma^{-1}}{3(1-\gamma^{2})^{2}q_{+}q_{-}}+\frac{4(q_{\gamma}-\gamma)}{(1-\gamma^{2})^{2}}\right]\|\boldsymbol{\mu}\|^{-2},$
where the last inequality is due to the condition on $\tilde{\varepsilon}$ of the regime (iii).
Finally, the desired conclusion holds by noting the following inequality
$\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}\geq\begin{cases}(1+\lambda)^{-1}\frac{n}{R}&\text{if $n\|\boldsymbol{\mu}\|^{2}\leq\lambda R$},\\ (1+\lambda^{-1})^{-1}\|\boldsymbol{\mu}\|^{-2}&\text{if $n\|\boldsymbol{\mu}\|^{2}\geq\lambda R$}.\end{cases}$
We now turn to a lower bound of $\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$.
Lemma B.9.
Suppose event $\bigcap_{i=1,2,3}\tilde{E}_{i}$ holds with $\tilde{\varepsilon}\leq\tfrac{q_{\gamma}}{2}$ and further assume one of the following conditions holds:
  1. (i)
    $q_{+}q_{-}=0$, $n\|\boldsymbol{\mu}\|^{2}<\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}}R$, $\tilde{\varepsilon}\leq\min\{\frac{\gamma^{3}}{4\sqrt{n}},\frac{\gamma^{5}}{16(1+\gamma)}\}$, and $\tilde{\varepsilon}_{2}\leq\frac{3\gamma^{2}}{16(1+\gamma)}\frac{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}{R}$,
  2. (ii)
    $q_{+}q_{-}\neq 0$, $n\|\boldsymbol{\mu}\|^{2}\leq\lambda R$, and $\tilde{\varepsilon}\leq\frac{q_{\gamma}^{2}}{4\big{\{}\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\big{\}}^{2}}\min\{\frac{1}{\sqrt{n}},\frac{q_{\gamma}}{4}\}$, $\tilde{\varepsilon}_{2}\leq\frac{3q_{\gamma}^{2}}{16\big{\{}\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\big{\}}^{2}}\frac{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}{R}$ for some $\lambda\leq\frac{q_{\gamma}}{\gamma(1-\gamma)(1+\tilde{\varepsilon}_{3})}$,
  3. (iii)
    $q_{\gamma}>\gamma$, $n\|\boldsymbol{\mu}\|^{2}\geq\lambda R$, $\tilde{\varepsilon}\leq\min\left\{\frac{1-\tilde{\varepsilon}_{3}}{4\sqrt{n}},\frac{q_{\gamma}}{8}\right\}q_{\gamma}^{2}\tilde{C}(\gamma,q_{+},\lambda)\frac{R}{n\|\boldsymbol{\mu}\|^{2}}$, $\tilde{\varepsilon}_{2}\leq\frac{3q_{\gamma}^{2}}{8\sqrt{n}}\tilde{C}(\gamma,q_{+},\lambda)$ for some $\lambda\geq\tfrac{\gamma}{4}$, and $\tilde{\varepsilon}_{3}\leq\tfrac{1}{2}$.
Then, there exists a constant $\tilde{c}_{2}$ which depends only on $q_{+}$ and $\gamma$ such that we have
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle\geq\tilde{c}_{2}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
Proof.
By (56) and Lemma B.5, we have
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$ $\geq(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)(\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}+Z\boldsymbol{\mu})$ (65)
$\qquad-\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\|B^{-1}\boldsymbol{y}\|\|B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)(\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}+Z\boldsymbol{\mu})\|$
$\geq\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$
$\qquad-R^{-1}\|B^{-1}\boldsymbol{y}\|\|B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)Z\boldsymbol{\mu}\|$
$\qquad\qquad-\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}\|B^{-1}\boldsymbol{y}\|\|B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)(\|\boldsymbol{\mu}\|^{2}\boldsymbol{y}+Z\boldsymbol{\mu})\|$
$\geq\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$
$\qquad-\frac{q_{\gamma}^{-1}\sqrt{n}\|Z\boldsymbol{\mu}\|}{R}-\frac{2q_{\gamma}^{-1}\tilde{\varepsilon}}{R}q_{\gamma}^{-1}\sqrt{n}(\sqrt{n}\|\boldsymbol{\mu}\|^{2}+\|Z\boldsymbol{\mu}\|)$
$\geq\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$
$\qquad-\frac{2n\|\boldsymbol{\mu}\|^{2}}{q_{\gamma}^{2}R}\tilde{\varepsilon}-\frac{q_{\gamma}+2\tilde{\varepsilon}}{3q_{\gamma}^{2}}\sqrt{n}\varepsilon_{2},$
where the second and third inequalities are due to $\|B^{-1}\|^{2}\leq q_{\gamma}^{-1}$ and $\|q_{+}B_{+}+q_{-}B_{-}\|\leq 1$, and the last inequality is due to the definition of $\tilde{\varepsilon}_{2}$.
We split in cases to bound the first term $\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$.
Case $q_{+}=1,q_{-}=0$: In this case, the first term becomes $\|\boldsymbol{\mu}\|^{2}(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}\boldsymbol{y}$. By (36), we have
$\|\boldsymbol{\mu}\|^{2}(B_{+}^{-1}\boldsymbol{y})^{\top}A^{-1}\boldsymbol{y}$ (66)
$=$ $\frac{\|\boldsymbol{\mu}\|^{2}}{R}\left\{\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)n_{+}\right.$
$\qquad\qquad\left.+\left(\gamma^{-1}-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)n_{-}\right\}$
$\geq$ $\frac{\|\boldsymbol{\mu}\|^{2}}{R}\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\frac{n_{+}+\gamma^{-1}n_{-}}{n}\right)(n_{+}+n_{-})$
$\geq$ $\left(1-\frac{n\|\boldsymbol{\mu}\|^{2}}{(1+\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}})R}\gamma^{-1}\right)\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$\geq$ $\left(1-\frac{\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}}}{1+\frac{\gamma}{1+(1+\gamma)\tilde{\varepsilon}_{3}}}\gamma^{-1}\right)\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$=$ $\frac{\gamma+(1+\gamma)\tilde{\varepsilon}_{3}}{(1+\gamma)(1+\tilde{\varepsilon}_{3})}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$\geq$ $\frac{\gamma}{2(1+\gamma)}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}.$
Thus, by (65) and noting $\tilde{\varepsilon}\leq\frac{q_{\gamma}}{2}$, we have for the regime (i)
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle\geq\left(\frac{\gamma}{2(1+\gamma)}-\frac{2\tilde{\varepsilon}}{\gamma^{4}}-\frac{2}{3\gamma^{2}}\frac{R}{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}\tilde{\varepsilon}_{2}\right)\frac{n\|\boldsymbol{\mu}\|^{2}}{R}\geq\frac{\gamma}{4(1+\gamma)}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}.$
Case $q_{+}=0,q_{-}=1$: In this case, the first term becomes $\|\boldsymbol{\mu}\|^{2}(B_{-}^{-1}\boldsymbol{y})^{\top}A^{-1}\boldsymbol{y}$. Similarly, under the regime (i), (39) implies the same lower bound as the case $q_{+}=1,q_{-}=0$.
Case $q_{+}q_{-}\neq 0$: By (45) and $B^{-1}(q_{+}B_{+}\boldsymbol{y}+q_{-}B_{-}\boldsymbol{y})=\frac{q_{+}+\gamma q_{-}}{\sqrt{q_{+}+\gamma^{2}q_{-}}}\boldsymbol{1}_{+}-\frac{q_{-}+\gamma q_{+}}{\sqrt{q_{-}+\gamma^{2}q_{+}}}\boldsymbol{1}_{-}$, we have
$\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$ (67)
$=$ $\|\boldsymbol{\mu}\|^{2}\{d(q_{+}+\gamma^{2}q_{-})(q_{-}+\gamma^{2}q_{+})\}^{-1}$
×[$\,\times\left[\left\{\left(q_{-}+\gamma^{2}q_{+}-\gamma\right)n_{-}\|\boldsymbol{\mu}\|^{2}+(q_{-}+\gamma^{2}q_{+})R\right\}(q_{+}+\gamma q_{-})n_{+}\right.$
+$\,\left.+\left\{\left(q_{+}+\gamma^{2}q_{-}-\gamma\right)n_{+}\|\boldsymbol{\mu}\|^{2}+(q_{+}+\gamma^{2}q_{-})R\right\}(q_{-}+\gamma q_{+})n_{-}\right].$
Under the regime (ii), (48), (49), and (67) imply
$\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$ (68)
$\geq$ $\|\boldsymbol{\mu}\|^{2}\frac{\frac{q_{\gamma}R}{2}\{(q_{+}+\gamma q_{-})n_{+}+(q_{-}+\gamma q_{+})n_{-}\}}{q_{\gamma}\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}R^{2}}$
$\geq$ $\|\boldsymbol{\mu}\|^{2}\frac{\frac{q_{\gamma}^{2}R}{2}(n_{+}+n_{-})}{q_{\gamma}\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}R^{2}}$
where the second inequality follows from $\min\{q_{+}+\gamma q_{-},q_{-}+\gamma q_{-}\}\geq q_{\gamma}$
With this, (65), and $\tilde{\varepsilon}\leq\frac{q_{\gamma}}{2}$, we have under regime (ii)
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$ $\geq\frac{q_{\gamma}}{2\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}-\frac{2n\|\boldsymbol{\mu}\|^{2}}{q_{\gamma}^{2}R}\tilde{\varepsilon}-\frac{q_{\gamma}+2\tilde{\varepsilon}}{3q_{\gamma}^{2}}\sqrt{n}\varepsilon_{2}$ (69)
=$=\left\{\frac{q_{\gamma}}{2\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}}-\frac{2}{q_{\gamma}^{2}}\tilde{\varepsilon}-\frac{q_{\gamma}+\tilde{\varepsilon}}{3q_{\gamma}^{2}}\frac{R}{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}\tilde{\varepsilon}_{2}\right\}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$\geq\left\{\frac{q_{\gamma}}{2\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}}-\frac{2}{q_{\gamma}^{2}}\tilde{\varepsilon}-\frac{2}{3q_{\gamma}}\frac{R}{\sqrt{n}\|\boldsymbol{\mu}\|^{2}}\tilde{\varepsilon}_{2}\right\}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$\geq\frac{q_{\gamma}}{4\left(\frac{(1+\tilde{\varepsilon}_{3})\lambda}{2}+1\right)^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$\geq\frac{q_{\gamma}}{4\left(1+\lambda\right)^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{R}$
$\geq\frac{q_{\gamma}}{4\left\{1+\frac{q_{\gamma}}{\gamma(1-\gamma)}\right\}^{2}}\frac{n\|\boldsymbol{\mu}\|^{2}}{R},$
where the last two lines are due to $\tilde{\varepsilon}\leq 1$ and $\lambda\leq\frac{q_{\gamma}}{\gamma(1-\gamma)(1+\tilde{\varepsilon}_{3})}$.
Under regime (iii), the definition of $\tilde{E}_{3}$, (51), and (67) imply
$\|\boldsymbol{\mu}\|^{2}(B^{-1}\boldsymbol{y})^{\top}A^{-1}B^{-1}\left(q_{+}B_{+}+q_{-}B_{-}\right)\boldsymbol{y}$ (70)
$\geq$ $\|\boldsymbol{\mu}\|^{2}\left\{\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\right\}^{-1}(n\|\boldsymbol{\mu}\|^{2})^{-2}$
×$\quad\times\left\{(q_{\gamma}-\gamma)(n_{+}+n_{-})\|\boldsymbol{\mu}\|^{2}+2q_{\gamma}R\right\}q_{\gamma}(n_{+}+n_{-})$
$\geq$ $\|\boldsymbol{\mu}\|^{2}\left\{\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}\right\}^{-1}(n\|\boldsymbol{\mu}\|^{2})^{-2}\cdot(q_{\gamma}-\gamma)q_{\gamma}n^{2}\|\boldsymbol{\mu}\|^{2}$
$=$ $\frac{q_{\gamma}(q_{\gamma}-\gamma)}{\frac{(1-\gamma^{2})^{2}}{16}+q_{\gamma}(1+\lambda^{-1})\lambda^{-1}}=q_{\gamma}\tilde{C}(\gamma,q_{+},\lambda).$
Thus, by (65) and noting $\tilde{\varepsilon}\leq\frac{q_{\gamma}}{2}$, we have under regime (iii),
$\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle$ $\geq q_{\gamma}\tilde{C}(\gamma,q_{+},\lambda)-\frac{2n\|\boldsymbol{\mu}\|^{2}}{q_{\gamma}^{2}R}\tilde{\varepsilon}-\frac{2}{3q_{\gamma}}\sqrt{n}\varepsilon_{2}$ (71)
$\geq\frac{q_{\gamma}\tilde{C}(\gamma,q_{+},\lambda)}{2}$
$\geq\frac{q_{\gamma}\tilde{C}(\gamma,q_{+},\lambda=\tfrac{\gamma}{4})}{2},$
where the last line is due to $\lambda\geq\tfrac{\gamma}{4}$.
Finally, the desired conclusion holds by noting the following inequallity
$\max\left\{\frac{n\|\boldsymbol{\mu}\|^{2}}{R},1\right\}\geq\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
Lemma B.10.
Consider model (M). If $\boldsymbol{w}\in\mathbb{R}^{p}$ is such that $\langle\boldsymbol{w},\boldsymbol{\mu}\rangle>0$, Then, we have
$\mathbb{P}(\langle\boldsymbol{w},y\boldsymbol{x}\rangle\leq 0)\leq\exp\left\{-c\left(\frac{\langle\boldsymbol{w},\boldsymbol{\mu}\rangle}{\|\boldsymbol{w}\|\|\boldsymbol{z}\|_{\psi_{2}}}\right)^{2}\right\},$
where $c$ is a universal constant and $\|\boldsymbol{z}\|_{\psi_{2}}$ is the sub-Gaussian norm of $\boldsymbol{z}$.
Proof.
Since $y\boldsymbol{x}=\boldsymbol{\mu}+y\boldsymbol{z}$, we have
$\mathbb{P}(\langle\boldsymbol{w},y\boldsymbol{x}\rangle\leq 0)$ $=\mathbb{P}(y\langle\boldsymbol{w},\boldsymbol{z}\rangle\leq-\langle\boldsymbol{w},\boldsymbol{\mu}\rangle)$
$=\frac{1}{2}\mathbb{P}(\langle\boldsymbol{w},\boldsymbol{z}\rangle\leq-\langle\boldsymbol{w},\boldsymbol{\mu}\rangle)+\frac{1}{2}\mathbb{P}(-\langle\boldsymbol{w},\boldsymbol{z}\rangle\leq-\langle\boldsymbol{w},\boldsymbol{\mu}\rangle)$
$=\mathbb{P}(|\langle\boldsymbol{w},\boldsymbol{z}\rangle|>\langle\boldsymbol{w},\boldsymbol{\mu}\rangle).$
Then the desired conclusion follows from the basic property of sub-Gaussian distributions (see Proposition 2.5.2, Vershynin [2018]). ∎
Theorem B.11 (Detailed version of Theorem 5.1).
Under the assumptions of Lemma B.9, we have
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(\boldsymbol{x};\hat{W})<0)\leq\exp\left(-c\frac{n\|\boldsymbol{\mu}\|^{4}}{\|\boldsymbol{z}\|_{\psi_{2}}^{2}\{n\|\boldsymbol{\mu}\|^{2}+R\}}\right),$
where $c$ is a constant which depends only on $q_{+}$ and $\gamma$.
Proof.
By Lemma B.8 and B.9, we have
$\left(\frac{\langle\boldsymbol{\bar{w}},\boldsymbol{\mu}\rangle}{\|\boldsymbol{\bar{w}}\|}\right)^{2}\geq\frac{\left(\tilde{c}_{2}\frac{n\|\boldsymbol{\mu}\|^{2}}{n\|\boldsymbol{\mu}\|^{2}+R}\right)^{2}}{\tilde{c}_{1}\frac{n}{n\|\boldsymbol{\mu}\|^{2}+R}}=\frac{\tilde{c}_{2}^{2}}{\tilde{c}_{1}}\frac{n\|\boldsymbol{\mu}\|^{4}}{n\|\boldsymbol{\mu}\|^{2}+R}.$
Then, the desired conclusion holds from Lemma B.10. ∎
Appendix C Proof of Theorem 6.1 and 6.2
First we establish connection between events $E(\theta_{1},\theta_{2})$ and $\tilde{E}_{1}(\tilde{\varepsilon}_{1},R),\tilde{E}_{2}(\varepsilon_{2},R)$.
Lemma C.1.
Suppose $\tilde{E}_{1}(\tilde{\varepsilon}_{1},R)\cap\tilde{E}_{2}(\varepsilon_{2},R)$ hold with $\tilde{\varepsilon}_{1}\leq\tfrac{1}{2}$, then event $E(\theta_{1},\theta_{2})$ holds with
$\theta_{1}=\frac{\tilde{\varepsilon}_{1}}{2}\quad\text{and}\quad\theta_{2}=\frac{\tilde{\varepsilon}_{2}\sqrt{R}}{2\|\boldsymbol{\mu}\|}.$
Proof.
By $\tilde{E}_{1}(\tilde{\varepsilon}_{1},R)$, we have $|\|\boldsymbol{z}_{i}\|^{2}-R|\leq\tfrac{\tilde{\varepsilon}_{1}}{3}R$ and $|\langle\boldsymbol{z}_{i},\boldsymbol{z}_{k}\rangle|\leq\tfrac{\tilde{\varepsilon}_{1}}{3}R$ for any $i\neq k$. Then, we have
$\Big{|}\Big{\langle}\frac{\boldsymbol{z}_{i}}{\|\boldsymbol{z}_{i}\|},\frac{\boldsymbol{z}_{k}}{\|\boldsymbol{z}_{k}\|}\Big{\rangle}\Big{|}$ $\leq\frac{\frac{\tilde{\varepsilon}_{1}}{3}R}{(1-\frac{\tilde{\varepsilon}_{1}}{3})R}$
$\leq\frac{\tilde{\varepsilon}_{1}}{2},$
where the last line follows from the assumption $\tilde{\varepsilon}_{1}\leq\tfrac{1}{2}$.
Similarly, by $\tilde{E}_{1}(\tilde{\varepsilon}_{1},R)\cap\tilde{E}_{2}(\varepsilon_{2},R)$ , we have
$\Big{|}\Big{\langle}\frac{\boldsymbol{z}_{i}}{\|\boldsymbol{z}_{i}\|},\frac{\boldsymbol{\mu}}{\|\boldsymbol{\mu}\|}\Big{\rangle}\Big{|}\leq\frac{\frac{\tilde{\varepsilon}_{2}}{3}R}{\sqrt{(1-\frac{\tilde{\varepsilon}_{1}}{3})R}\|\boldsymbol{\mu}\|}\leq\frac{\tilde{\varepsilon}_{2}\sqrt{R}}{2\|\boldsymbol{\mu}\|},$
where the last inequality is due to $\tilde{\varepsilon}_{1}\leq\tfrac{1}{2}$. ∎
For model (sG), we have
Lemma C.2 (Lemma S4.1, S4.2, Hashimoto et al. [2025]).
Consider model (sG). There exist universal constants $C_{i},i=1,2$ such that for any $\delta\in(0,\tfrac{1}{2})$, event $\tilde{E}_{1}\cap\tilde{E}_{2}$ holds with probability at least $1-2\delta$, where
$R$ $={\rm tr}(\Sigma),$
$\tilde{\varepsilon}_{1}$ $:=C_{1}\max\left\{\sqrt{n\log\left(\tfrac{1}{\delta}\right)}\|\Sigma\|_{F},n\log\left(\tfrac{1}{\delta}\right)\|\Sigma\|\right\}/{\rm tr}(\Sigma)$
$\tilde{\varepsilon}_{2}$ $:=C_{2}\sqrt{n\log\left(\tfrac{1}{\delta}\right)}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|/{\rm tr}(\Sigma).$
Lemma C.3.
There exists a universal constant $C_{3}$ such that we have with probability at least $1-\delta$
$|n_{+}-n_{-}|\leq C_{3}\sqrt{n\log(\tfrac{1}{\delta})}.$
Proof.
By independence of $y_{i}$ and $y_{i}\in\{\pm 1\}$, Hoeffding’s inequality implies that
$\left|\sum_{i}y_{i}\right|\leq C_{3}\sqrt{n\log(\tfrac{1}{\delta})}$
holds with probability at least $1-\delta$, where $C_{3}$ is a universal constant.
The desired result follows by noting that $\sum_{i}y_{i}=n_{+}-n_{-}$. ∎
Corollary C.4.
Event $\tilde{E}_{3}$ holds with $\tilde{\varepsilon}_{3}:=C_{3}\sqrt{\frac{\log(\frac{1}{\delta})}{n}}$ with probability at least $1-\delta$.
With these, we can rewrite assumptions of our main theorems, namely Theorem 4.4, 4.9, and B.11, using parameters defining the model (sG).
From Theorem 4.4 and 4.9, we have the following result on (sG).
Theorem C.5.
Consider model (sG). Suppose
${\rm tr}(\Sigma)\geq\max\left\{2C_{1}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma\|_{F},2C_{1}n\log(\tfrac{1}{\delta})\|\Sigma\|,C_{2}^{2}n\log(\tfrac{1}{\delta})\frac{\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|^{2}}{\|\boldsymbol{\mu}\|^{2}}\right\},$ (72)
and further suppose one of the following conditions:
  1. (i)
    $\|\boldsymbol{\mu}\|^{2}\geq 2.18\max\big{\{}C_{1}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma\|_{F},C_{1}n\log(\tfrac{1}{\delta})\|\Sigma\|,C_{2}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\big{\}}$,
    $\alpha\leq\left[1.2\{n\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)\}\right]^{-1}$, and $\sigma\leq 0.2\gamma\alpha\sqrt{\frac{\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)}{m}}$,
  2. (ii)
    ${\rm tr}(\Sigma)\geq\frac{67.2e}{\gamma^{2}}n\|\boldsymbol{\mu}\|^{2}$,
    ${\rm tr}(\Sigma)\geq\frac{453e}{\gamma^{3}}n\max\big{\{}C_{1}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma\|_{F},C_{1}n\log(\tfrac{1}{\delta})\|\Sigma\|,C_{2}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\big{\}}$,
    $\alpha\leq\{8{\rm tr}(\Sigma)\}^{-1}$, and $\sigma\leq 0.1\gamma\alpha\sqrt{\frac{{\rm tr}(\Sigma)}{m}}$.
Then, we have with probability at least $1-2\delta$, the gradient descent iterate (2) keeps all the neurons activated for $t\geq 1$, satisfies $\mathcal{L}(W^{(t)})=O(t^{-1})$, and converges in the direction given by Theorem 4.1.
Proof.
By Lemma C.1, C.2, and the assumption (72), all the conclusions of Lemma C.1 and C.2 hold with probability at least $1-2\delta$ with
$R$ $={\rm tr}(\Sigma),$
$\theta_{1}$ $=\frac{\tilde{\varepsilon}_{1}}{2}=\frac{C_{1}\max\left\{\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma\|_{F},n\log(\tfrac{1}{\delta})\|\Sigma\|\right\}}{2{\rm tr}(\Sigma)}\leq\frac{1}{4},$
$\theta_{2}$ $=\frac{\tilde{\varepsilon}_{2}\sqrt{R}}{2\|\boldsymbol{\mu}\|}=\frac{C_{2}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|}{2\sqrt{{\rm tr}(\Sigma)}\|\boldsymbol{\mu}\|}\leq\frac{1}{2}.$
Then, we have $0.91\sqrt{{\rm tr}(\Sigma)}\leq R_{min}(\boldsymbol{z})\leq R_{max}(\boldsymbol{z})\leq 1.09\sqrt{{\rm tr}(\Sigma)}$.
From these, we have
$2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ (73)
$\leq$ $1.09C_{2}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|$
$\qquad\qquad+0.6C_{1}\max\left\{\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma\|_{F},n\log(\tfrac{1}{\delta})\|\Sigma\|\right\}$
$\leq$ $2.18\max\{C_{1}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma\|_{F},C_{1}n\log(\tfrac{1}{\delta})\|\Sigma\|,C_{2}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\}.$
Under condition (i), the desired conclusions follow from Theorem 4.4.
First, it is immediate from (73) to see $\|\boldsymbol{\mu}\|^{2}\geq 2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$ holds under condition (i).
Furthermore, condition (i) also implies that Assumption 4.3 is satisfied since
$\alpha\{n\|\boldsymbol{\mu}\|^{2}+R_{max}^{2}(\boldsymbol{z})\}<1.2\alpha\{n\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)\}\leq 1.$
We note that $\exp(\rho)<1.26$ holds under condition (i) since
$\rho$ $\leq\sigma\sqrt{m\cdot\frac{3}{2}\cdot 1.2\{\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)\}}$
$\leq 1.35\sigma\sqrt{m\{\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)\}}$
$\leq 0.27\gamma\alpha\{\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)\}$
$\leq\frac{0.27}{1.2}\gamma\leq 0.225.$
where the first inequality follows from the definition of $\rho$ and $\theta_{2}\leq\tfrac{1}{2}$ and $R_{max}(\boldsymbol{z})\leq 1.2\sqrt{{\rm tr}(\Sigma)}$, the remaining inequalities follow from condition (i).
Noting $\exp(\rho)<1.26$ and $\rho\leq 0.27\gamma\alpha\{\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)\}$, we also see that Assumption 4.2 is satisfied since
$\rho\exp(\rho)$ $<1.26\rho$
$\leq 1.26\cdot 0.27\gamma\alpha\{\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)\}$
$\leq\frac{1.26\cdot 0.27}{0.8}\gamma\alpha\{\|\boldsymbol{\mu}\|^{2}+R_{min}^{2}(\boldsymbol{z})\}$
$\leq\gamma(1-\theta_{2})\alpha\{\|\boldsymbol{\mu}\|^{2}+R_{min}^{2}(\boldsymbol{z})\}.$
Since all the assumptions of Theorem 4.4 are satisfied, the desired conclusions hold with probability at least $1-2\delta$ under condition (i).
Under condition (ii), the desired conclusions follow from Theorem 4.9. To see this, first note that $\tilde{\varepsilon}_{1}\leq\tfrac{1}{2}$ implies that $\frac{R_{max}^{2}(\boldsymbol{z})}{R_{min}^{2}(\boldsymbol{z})}\leq\frac{1+\frac{\tilde{\varepsilon}_{1}}{3}}{1-\frac{\tilde{\varepsilon}_{1}}{3}}\leq\frac{7}{5}$. So, we have $C_{w}$ in Theorem 4.9 with $C_{w}\leq\tfrac{168e}{5\gamma^{2}}$. So, by setting $\varepsilon_{i}=\tfrac{5\gamma^{2}}{368e},i=1,3$, $C_{w}\varepsilon_{i}\leq\tfrac{1}{2}$ is ensured. Moreover, by setting $\varepsilon_{2}=\tfrac{1}{2}$, $\varepsilon_{1}+\varepsilon_{2}+\varepsilon_{3}<1$ is ensured.
We note that $\exp(\rho)<\frac{10}{9}$ holds under condition (ii) since
$\rho$ $\leq 1.35\sigma\sqrt{m\{\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)\}}$
$\leq 1.35\sigma\sqrt{m(1.0055){\rm tr}(\Sigma)}$
$\leq 1.36\sigma\sqrt{m{\rm tr}(\Sigma)}$
$\leq 1.36\cdot 0.1\gamma\alpha{\rm tr}(\Sigma)$
$\leq\frac{1.36\cdot 0.1}{8}\gamma\leq 0.017<\log(\tfrac{10}{9}).$
By noting $\rho\leq 0.136\gamma\alpha{\rm tr}(\Sigma)$ and $\exp(\rho)<\frac{10}{9}$, condition (ii) implies that Assumption 4.6 is satisfied since
$\rho\exp(\rho)$ $\leq\tfrac{10}{9}\cdot 0.136\gamma\alpha{\rm tr}(\Sigma)$
$\leq\tfrac{10}{9}\cdot 0.136\cdot 0.8^{-1}\gamma\alpha R_{min}^{2}(\boldsymbol{z})$
$\leq 0.19\gamma\alpha R_{min}^{2}(\boldsymbol{z})$
$\leq\varepsilon_{2}\gamma(1-\theta_{2})\alpha R_{min}^{2}(\boldsymbol{z}).$
Similarly, by noting $\exp(\rho)<\frac{10}{9}$, condition (ii) also implies that Assumption 4.8 is satisfied since
$6\alpha R_{min}^{2}(\boldsymbol{z})\exp(\rho)$ $<7.2\cdot\tfrac{10}{9}\alpha R_{min}^{2}(\boldsymbol{z})$
$=8\alpha{\rm tr}(\Sigma)\leq 1.$
By (73), condition (ii) implies that Assumption 4.5 is satisfied since
$2\theta_{2}\|\boldsymbol{\mu}\|R_{max}(\boldsymbol{z})+\theta_{1}R_{max}^{2}(\boldsymbol{z})$
$\leq$ $2.18\max\{C_{1}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma\|_{F},C_{1}n\log(\tfrac{1}{\delta})\|\Sigma\|,C_{2}\sqrt{n\log(\tfrac{1}{\delta})}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\}$
$\leq$ $\frac{\gamma^{3}{\rm tr}(\Sigma)}{453en}$
$\leq$ $\frac{\frac{5\gamma^{2}}{336e}\cdot\frac{1}{2}\cdot 0.8\gamma{\rm tr}(\Sigma)}{n(10/9)^{2}}$
$\leq$ $\frac{\varepsilon_{1}\gamma(1-\theta_{2})R_{min}^{2}(\boldsymbol{z})}{n\exp(2\rho)}.$
Since all the assumptions of Theorem 4.9 are satisfied, the desired conclusions hold with with probability at least $1-2\delta$ under condition (ii). ∎
Application of Theorem B.11 on the model (sG) is given below:
Theorem C.6.
Consider model (sG). Suppose $n\geq 4C_{3}^{2}\log(\frac{1}{\delta})$ and one of the following holds:
  1. (i)
    $q_{+}q_{-}=0$, $n\geq\frac{16(1+\gamma)^{2}}{\gamma^{4}}$, $n\|\boldsymbol{\mu}\|^{2}<\frac{\gamma}{1+(1+\gamma)/2}{\rm tr}(\Sigma)$,
    ${\rm tr}(\Sigma)$ $\geq 4\gamma^{-3}n\sqrt{\log(\tfrac{1}{\delta})}\max\left\{C_{1}\|\Sigma\|_{F},C_{1}\|\Sigma\|\sqrt{n\log(\tfrac{1}{\delta})},C_{2}\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\right\},\text{and}$
    $\|\boldsymbol{\mu}\|^{2}$ $\geq\frac{16C_{2}(1+\gamma)}{3\gamma^{2}}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\sqrt{\log(\tfrac{1}{\delta})},$
  2. (ii)
    $q_{+}q_{-}\neq 0$, $n\geq\frac{16}{q_{\gamma}^{2}}$, $n\|\boldsymbol{\mu}\|^{2}\leq\lambda{\rm tr}(\Sigma)$,
    ${\rm tr}(\Sigma)$ $\geq\frac{4(\frac{3\lambda}{4}+1)^{2}}{q_{\gamma}^{2}}n\sqrt{\log(\tfrac{1}{\delta})}\max\left\{C_{1}\|\Sigma\|_{F},C_{1}\|\Sigma\|\sqrt{n\log(\tfrac{1}{\delta})},C_{2}\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\right\},\text{and}$
    $\|\boldsymbol{\mu}\|^{2}$ $\geq\frac{16C_{2}(\frac{3\lambda}{4}+1)^{2}}{3q_{\gamma}^{2}}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\sqrt{\log(\tfrac{1}{\delta})},$
    for some $\lambda\leq\frac{2q_{\gamma}}{3\gamma(1-\gamma)}$,
  3. (iii)
    $q_{\gamma}>\gamma$, $n\geq\frac{4}{q_{\gamma}^{2}}$, $n\|\boldsymbol{\mu}\|^{2}\geq\lambda{\rm tr}(\Sigma)$,
    ${\rm tr}(\Sigma)$ $\geq\frac{2\sqrt{2}}{q_{\gamma}\sqrt{\tilde{C}(\gamma,q_{+},\lambda)}}n\|\boldsymbol{\mu}\|\{\log(\tfrac{1}{\delta})\}^{1/4}$
    ×$\qquad\times\sqrt{\max\left\{C_{1}\|\Sigma\|_{F},C_{1}\|\Sigma\|\sqrt{n\log(\tfrac{1}{\delta})},C_{2}\sqrt{n}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\right\}},\text{and}$
    ${\rm tr}(\Sigma)$ $\geq\frac{C_{2}}{3q_{\gamma}^{2}\tilde{C}(\gamma,q_{+},\lambda)}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|\sqrt{n\log(\tfrac{1}{\delta})}$
    for some $\lambda\geq\tfrac{\gamma}{4}$.
Then, we have with probability at least $1-3\delta$
$\mathbb{P}_{(\boldsymbol{x},y)}(yf(\boldsymbol{x};\hat{W})<0)\leq\exp\left(-c\frac{n\|\boldsymbol{\mu}\|^{4}}{L^{2}\|\Sigma\|\{n\|\boldsymbol{\mu}\|^{2}+{\rm tr}(\Sigma)\}}\right),$
where $c$ is a constant which depends only on $\gamma$ and $q_{+}$.
Proof.
By Lemma C.2 and Corollary C.4, event $\bigcap_{i=1,2,3}\tilde{E}_{i}$ holds with probability at least $1-3\delta$ with
$R$ $={\rm tr}(\Sigma),$
$\tilde{\varepsilon}_{1}$ $=C_{1}\max\left\{\sqrt{n\log\left(\tfrac{1}{\delta}\right)}\|\Sigma\|_{F},n\log\left(\tfrac{1}{\delta}\right)\|\Sigma\|\right\}/{\rm tr}(\Sigma)$
$\tilde{\varepsilon}_{2}$ $=C_{2}\sqrt{n\log\left(\tfrac{1}{\delta}\right)}\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|/{\rm tr}(\Sigma),$
$\tilde{\varepsilon}_{3}$ $=C_{3}\sqrt{\frac{\log(\tfrac{1}{\delta})}{n}}.$
Note that the assumption $n\geq 4C_{3}^{2}\log(\tfrac{1}{\delta})$ implies $\tilde{\varepsilon}_{3}\leq\tfrac{1}{2}$.
Then, the desired conclusion follows by simply rewriting assumptions of Theorem B.11. ∎
We can now see that Theorem 6.1 and 6.2 follow from Theorem C.5 and C.6 by ensuring both assumptions are satisfied. Specifically, Theorem C.5 under condition (i) and Theorem C.6 under condition (iii) imply Theorem 6.1. Similarly, Theorem C.5 under condition (i) and Theorem C.6 under condition (i) and (ii) imply Theorem 6.2 under condition (ii). Lastly, Theorem C.5 under condition (ii) and Theorem C.6 under condition (i) and (ii) imply Theorem 6.2 under condition (ii)